### 15 Mar 2026 v1

# Learning-to-Defer with Expert-Conditioned Advice

Abstract
Learning-to-Defer routes each input to the expert that minimizes expected cost, but it assumes that the information available to every expert is fixed at decision time. Many modern systems violate this assumption: after selecting an expert, one may also choose what additional information that expert should receive, such as retrieved documents, tool outputs, or escalation context. We study this problem and call it Learning-to-Defer with advice. We show that a broad family of natural separated surrogates, which learn routing and advice with distinct heads, are inconsistent even in the smallest non-trivial setting. We then introduce an augmented surrogate that operates on the composite expert–advice action space and prove an $\mathcal{H}$-consistency guarantee together with an excess-risk transfer bound, yielding recovery of the Bayes-optimal policy in the limit. Experiments on tabular, LLMs, and multi-modal tasks show that the resulting method improves over standard Learning-to-Defer while adapting its advice-acquisition behavior to the cost regime.
1 Introduction
Modern decision systems rarely act on a fixed snapshot of information. A retrieval-augmented model may be chosen to generate an answer and then be provided with specific retrieved documents (Guu et al., 2020; Lewis et al., 2020). A tool-augmented agent may be assigned a task and then granted access to a calculator, a search engine, or a code interpreter (Schick et al., 2023). In each of these settings, two decisions are made in sequence: who should handle the instance, and what additional information, if any, should be acquired to help them. The value of that additional information is not known when the decision is made; it is revealed only after the commitment. What, then, should the learner optimize: the sequential control flow, or the final executed decision?
Learning-to-Defer provides a principled framework for the first of these two decisions. Given an input $x$, the system routes each instance to the expert best suited to handle it, balancing predictive performance against expert-specific costs (Madras et al., 2018; Mozannar and Sontag, 2020; Mao et al., 2023a; Montreuil et al., 2026a). This framework has produced strong theoretical foundations and practical successes. Yet it operates under a fundamental assumption: the information available to every expert is fixed at decision time. The system chooses who acts, but never what they get to see. The entire question of whether and what to acquire is outside the model. This misses an important class of systems in which the value of extra information depends on the expert who will consume it, and can even change which expert is optimal. Recent work has begun to study the binary ask/not-ask case of querying human feedback for a single enriched predictor (Pugnana et al., 2025); our focus here is the more general routed setting in which the value of advice depends on the expert that will use it. In this richer setting, the central statistical question is Bayes consistency: can one learn a policy that recovers the Bayes-optimal executed expert-advice decision, or equivalently the best expert under its best advice choice?
This coupling makes the problem qualitatively different from standard learning to defer. A natural response is to keep the sequential structure and learn routing and advice with separate heads. But does such a decomposition preserve the Bayes rule of the underlying protocol? We show that, in general, it does not. Even in the simplest setting — two experts and a single binary advice — the separated surrogate is inconsistent for the true loss. The issue is not optimization, but statistical alignment: the protocol executes an expert–advice pair, while the separated surrogate compares the wrong objects.
We call this framework Learning-to-Defer with advice. It generalizes classical Learning-to-Defer by introducing expert-conditioned advice acquisition as a first-class component of the deferral decision. Can one still obtain a learning problem that is both statistically correct and simple to optimize, despite the negative result above? We show that the answer is yes. We introduce a surrogate with an explicit Bayes-consistency guarantee: its minimization is provably aligned with the Bayes-optimal executed expert-advice decision of the original protocol, that is, with the expert whose best advice yields the smallest conditional expected cost.
Contributions.
(i) we introduce Learning-to-Defer with advice, a formulation in which the system jointly selects an expert and an expert-conditional advice action, and we derive the corresponding Bayes-optimal policy, which compares experts at their best-advised cost (Section 4, especially Section 4.3 and Lemma 3); (ii) we prove that a broad family of natural separated router-query surrogates can be Fisher inconsistent for the true protocol loss, even in the simplest binary-advice setting (Section 4.4 and Theorem 6); (iii) we introduce an augmented surrogate and prove an $\mathcal{H}$-consistency guarantee together with an excess-risk transfer bound, thereby establishing Bayes consistency and showing that the learned policy recovers the Bayes-optimal expert-advice decision in the limit, equivalently the expert whose best advice yields the smallest conditional expected cost (Section 4.5, Lemma 7, Theorem 8, and Corollary 9); and (iv) we validate the framework on tabular, language, multi-modal, and synthetic experiments, where it improves over standard Learning-to-Defer, adapts its advice-acquisition rate to the cost regime, and empirically exhibits the failure mode predicted for separated surrogates (Section 5 and Appendix C.2).
2 Related Work
L2D extends selective prediction (Chow, 1970; Bartlett and Wegkamp, 2008; Cortes et al., 2016; Geifman and El-Yaniv, 2017; Cao et al., 2022; Cortes et al., 2024b) by allowing a learner to defer uncertain inputs to external experts (Madras et al., 2018; Mozannar and Sontag, 2020; Verma et al., 2022). A substantial line of work develops surrogate losses and statistical guarantees for this problem (Mozannar and Sontag, 2020; Verma et al., 2022; Cao et al., 2024; Mozannar et al., 2023; Mao et al., 2024b, 2025; Charusaie et al., 2022; Mao et al., 2024a; Wei et al., 2024; Montreuil et al., 2026a). The framework has also been extended to regression and multi-task settings and deployed in real systems (Mao et al., 2024c; Strong et al., 2024; Palomba et al., 2025; Montreuil et al., 2025b; Strong et al., 2025; Montreuil et al., 2026c), as well as to robust (Montreuil et al., 2025a, 2026d), online (Montreuil et al., 2026e, b), and budgeted or missing experts (DeSalvo et al., 2025; Nguyen et al., 2025) settings.
Recent work has also considered querying human feedback and feeding it to an enriched predictor rather than treating the human only as a deferred decision maker (Pugnana et al., 2025). Their formulation is a binary ask/not-ask problem: a selector chooses between a standard predictor and a single feedback-augmented predictor, with no routing decision over experts. On the theory side, they derive an optimal thresholding rule and establish realizable-consistency guarantees, but for our purposes the key limitation is that Bayes consistency is not established. Our setting is richer: rather than a binary ask/not-ask gate, the learner must choose an executed expert-advice pair, or equivalently the expert whose best advice yields the smallest conditional expected cost. This stronger coupling is exactly what gives rise to our negative result for separated surrogates and makes Bayes consistency the central theoretical requirement addressed in this paper.
3 Preliminaries
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space supporting a pair of random variables $(X,Y)$ with joint distribution $\mathcal{D}$ on $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is the input space and $\mathcal{Y}$ is an arbitrary measurable output space. We denote by $(x,y)$ a realization of $(X,Y)$.
Learning-to-Defer.
We consider a fixed collection of $J$ experts, indexed by $J]\coloneqq\{1,\dots,J\}$. Each expert $j\in[J]$ is associated with a measurable cost function $c_{j}:\mathcal{X}\times\mathcal{Y}\to[0,C]$, uniformly bounded by some $C<\infty$, that quantifies the expense of routing $(x,y)$ to expert $j$. This formulation is agnostic to the nature of the expert: an expert may be a predictive model, a language model system, a human annotator, or any other decision mechanism that induces measurable costs. We model expert prediction as $e_{j}:\mathcal{X}\to\mathcal{Y}$ and decompose the cost as $c_{j}(x,y)\;\coloneqq\;\psi\bigl(e_{j}(x),\,y\bigr)+\beta_{j}$, where $\psi:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a measurable task-specific loss (e.g., the 011 loss for classification) and $\beta_{j}\geq 0$ is an expert consultation cost (Madras et al., 2018; Mozannar and Sontag, 2020; Verma et al., 2022). We denote the tuple of deterministic expert predictors, when this special case is used, by $\mathbf{e}\coloneqq(e_{1},\dots,e_{J})$.
Given a collection of fixed, pre-trained experts, Learning-to-Defer seeks a routing policy that assigns each instance to the most cost-effective expert (Madras et al., 2018; Mozannar and Sontag, 2020; Narasimhan et al., 2022; Mao et al., 2023a, 2024c; Montreuil et al., 2025b, 2026a). Let $\mathcal{H}_{r}\subseteq\{s_{r}:\mathcal{X}\times[J]\to\mathbb{R}\}$ denote a hypothesis class of measurable scoring functions for routing. Each $s_{r}\in\mathcal{H}_{r}$ induces a measurable router $r:\mathcal{X}\to[J]$ via $r(x)=\operatorname*{arg\,max}_{j\in[J]}s_{r}(x,j)$; we write $\mathbf{s}_{r}(x)=(s_{r}(x,j))_{j\in[J]}$ for the full score vector. The deferral loss incurred by routing a realization $(x,y)$ through $r$ is
$\ell_{\mathrm{def}}(r;\,x,y,\mathbf{e})\;\coloneqq\;\sum_{j\in[J]}c_{j}(x,y)\,\mathbf{1}\{r(x)=j\},$ (1)
and the population risk is its expectation $\mathcal{E}_{\ell_{\mathrm{def}}}(r)\;\coloneqq\;\mathbb{E}_{(X,Y)\sim\mathcal{D}}\bigl[\ell_{\mathrm{def}}(r;\,X,Y,\mathbf{e})\bigr].$
$\mathcal{H}_{r}$-consistency.
The indicator functions in $\ell_{\mathrm{def}}$ make it discontinuous and non-differentiable, precluding direct gradient-based optimization (Zhang, 2002; Steinwart, 2007; Mozannar and Sontag, 2020). Standard practice replaces it with a convex surrogate $\Phi_{\mathrm{def}}$ that operates on the scoring function $s_{r}$ and is amenable to gradient-based methods. The surrogate deferral risk is $\mathcal{E}_{\Phi_{\mathrm{def}}}(s_{r})=\mathbb{E}_{(X,Y)\sim\mathcal{D}}[\Phi_{\mathrm{def}}(s_{r};\,X,Y,\mathbf{e})]$, with best-in-class value $\mathcal{E}_{\Phi_{\mathrm{def}}}^{\ast}(\mathcal{H}_{r})\coloneqq\inf_{s_{r}\in\mathcal{H}_{r}}\mathcal{E}_{\Phi_{\mathrm{def}}}(s_{r})$. To obtain population-level excess-risk control, Awasthi et al. (2022) introduced $\mathcal{H}$-consistency bounds. Applied to $\mathcal{H}_{r}$, the bound takes the following form.
Definition 1 ($\mathcal{H}_{r}$-consistency bound).
The surrogate $\Phi_{\mathrm{def}}$ satisfies an $\mathcal{H}_{r}$-consistency bound with respect to $\ell_{\mathrm{def}}$ if there exists a non-decreasing function $\Gamma:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $\Gamma(0)=0$ such that for all $s_{r}\in\mathcal{H}_{r}$,
$\mathcal{E}_{\ell_{\mathrm{def}}}(r)-\mathcal{E}_{\ell_{\mathrm{def}}}^{B}(\mathcal{H}_{r})+\mathcal{U}_{\ell_{\mathrm{def}}}(\mathcal{H}_{r})\;\leq\;\Gamma\!\Bigl(\mathcal{E}_{\Phi_{\mathrm{def}}}(s_{r})-\mathcal{E}_{\Phi_{\mathrm{def}}}^{\ast}(\mathcal{H}_{r})+\mathcal{U}_{\Phi_{\mathrm{def}}}(\mathcal{H}_{r})\Bigr),$ (2)
where $r$ is the router induced by $s_{r}$, and $\mathcal{U}_{\ell_{\mathrm{def}}}(\mathcal{H}_{r})=\mathcal{E}_{\ell_{\mathrm{def}}}^{B}(\mathcal{H}_{r})-\mathbb{E}_{X}\bigl[\inf_{s_{r}\in\mathcal{H}_{r}}\mathbb{E}_{Y\mid X}[\ell_{\mathrm{def}}(r;\,X,Y,\mathbf{e})]\bigr]$ is the minimizability gap.
When $\mathcal{H}_{r}$ is rich enough for these gaps to vanish, the bound recovers Bayes consistency (Steinwart, 2007; Awasthi et al., 2022). In the standard L2D setting, such bounds have been established for the comp-sum family of cross-entropy surrogates (Mao et al., 2023b, 2024a), guaranteeing that surrogate minimization over a rich enough class recovers the Bayes-optimal router. In the next section, we extend the routing problem to include advice acquisition and examine which surrogates preserve consistency in this richer setting.
4 Problem Formulation and Surrogate Design
The deferral framework of Section 3 assumes that every expert acts on the input $x$ alone. We now extend this setting by allowing the system to acquire additional information—advice—and provide it to the chosen expert before that expert acts. This introduces a genuine statistical complication: the policy must choose both the expert and the advice source from $x$ alone, while the realized advice is revealed only after that commitment. We therefore first formalize the protocol and its Bayes rule, and then study which surrogate constructions remain valid in this extended setting.
4.1 Advice Sources
Advice acquisition introduces two coupled difficulties. First, the value of a source is expert-dependent: the same retrieved passage or tool call may help one expert and be irrelevant, or even harmful, for another (Guu et al., 2020; Lewis et al., 2020; Schick et al., 2023). The choice of what to acquire therefore cannot be assessed independently of the choice of who receives it. Second, both decisions must be made from $x$ alone. The realized advice value is revealed only after the expert and advice source have already been selected. The difficulty is therefore not merely that the protocol is sequential, but that learning must target a hidden executed expert-advice pair rather than two independently observable decisions.
Formal setup.
We formalize this uncertainty by augmenting the data-generating process with $K$ advice variables. The triple $(X,A,Y)$ is jointly distributed, where $A=(A^{1},\dots,A^{K})$ and each $A^{k}$ takes values in a measurable space $\mathcal{A}_{k}$. Let $\mathcal{A}\coloneqq\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{K}$, and write $(x,a,y)$ for a realization. This model accommodates both deterministic advice, such as a database lookup, and stochastic advice, such as a language model output (Touvron et al., 2023; OpenAI et al., 2024). The action $k=0$ denotes the decision to acquire no advice, and $K]_{0}\coloneqq\{0,1,\dots,K\}$ is the resulting advice-action set.
Masking.
The protocol enforces that at most one source is revealed per instance. Fix a missing-advice symbol $\bot\notin\bigcup_{k=1}^{K}\mathcal{A}_{k}$ and define $\widetilde{\mathcal{A}}\coloneqq(\mathcal{A}_{1}\cup\{\bot\})\times\cdots\times(\mathcal{A}_{K}\cup\{\bot\})$. For each $k\in[K]_{0}$, the masking operator $m_{k}:\mathcal{A}\to\widetilde{\mathcal{A}}$ selects the $k$-th coordinate and masks all others:
$\bigl(m_{k}(a)\bigr)^{i}\;\coloneqq\;\begin{cases}\bot&\text{if }k=0,\\ a^{i}&\text{if }k\in[K]\text{ and }i=k,\\ \bot&\text{if }k\in[K]\text{ and }i\neq k,\end{cases}$ (3)
Thus every expert receives an element of $\widetilde{\mathcal{A}}$, but at most one coordinate contains a realized value. We write $\widetilde{a}^{(k)}\coloneqq m_{k}(a)$ and $\widetilde{A}^{(k)}\coloneqq m_{k}(A)$ for the masked advice.
4.2 The Deferral-Advice Loss
With advice sources and masking in place, we now define how the system acts and what it pays. The goal is to formalize the true loss of the protocol and identify the Bayes-optimal policy.
Costs.
In the standard setting, routing to expert $j$ incurs a cost $c_{j}(x,y)$ that captures both prediction error and consultation fee. With advice, the cost acquires a new dimension: the system may spend an acquisition fee $\gamma_{k}$ to provide advice to the expert, hoping that the additional information reduces the expert’s prediction error by more than it costs to obtain. For each pair $(j,k)\in[J]\times[K]_{0}$, let $c_{j,k}:\mathcal{X}\times\mathcal{A}\times\mathcal{Y}\to[0,C]$ be a measurable cost function, uniformly bounded by $C<\infty$.
$c_{j,k}(x,a,y)\;\coloneqq\;\underbrace{\psi\bigl(e_{j}(x,\widetilde{a}^{(k)}),\,y\bigr)}_{\text{task loss}}+\underbrace{\beta_{j}}_{\text{expert fee}}+\underbrace{\gamma_{k}}_{\text{advice fee}},$ (4)
where $e_{j}:\mathcal{X}\times\widetilde{\mathcal{A}}\to\mathcal{Y}$ extends the expert prediction of Section 3 to accept masked advice $\widetilde{a}^{(k)}$—reducing to $e_{j}(x)$ when k=0k=0—and $\beta_{j}\geq 0$ is the expert consultation cost, while $\gamma_{k}\geq 0$ is the advice acquisition cost with $\gamma_{0}=0$. The trade-off is explicit: acquiring advice $k$ incurs the fee $\gamma_{k}$ but may reduce the task loss $\psi$ if the advice is informative for expert $j$. Whether the trade-off is favorable depends jointly on the expert, the advice source, and the input.
Policy and true loss.
Formally, a deferral-advice policy is a pair of measurable maps $r:\mathcal{X}\to[J]$ and $q:\mathcal{X}\times[J]\to[K]_{0}$. Let $\mathcal{H}_{q}\subseteq\{s_{q}:\mathcal{X}\times[J]\times[K]_{0}\to\mathbb{R}\}$ be a hypothesis class of measurable query scoring functions; each $s_{q}\in\mathcal{H}_{q}$ induces a query function via $q(x,j)=\operatorname*{arg\,max}_{k\in[K]_{0}}s_{q}(x,j,k)$. For a fixed input $x$, let $\mathbf{S}_{q}(x)\in\mathbb{R}^{J\times(K+1)}$ denote the query-score matrix with entries $s_{q}(x,j,k)$. Equivalently, taking the argmax along each row yields one candidate advice index for each expert. Upon observing $x$, the system routes to expert $j=r(x)$ and selects advice index $k=q(x,j)$; the masked advice $\widetilde{A}^{(k)}$ is then revealed to expert $j$. The query function $q(x,\cdot)$ is defined for every expert, but only the executed value $q(x,r(x))$ enters the realized loss: the remaining values $q(x,j^{\prime})$ for $j^{\prime}\neq r(x)$ are counterfactual and invisible to the protocol. Appendix Algorithm 2 records this inference rule explicitly.
Definition 2 (True deferral-advice loss).
Given a policy $(r,q)$, the true deferral-advice loss on a realization $(x,a,y)$ is
$\ell_{\mathrm{def\text{-}adv}}(r,q;\,x,a,y,\mathbf{e})\;\coloneqq\;\sum_{j\in[J]}\sum_{k\in[K]_{0}}c_{j,k}(x,a,y)\,\mathbf{1}\{r(x)=j\}\,\mathbf{1}\{q(x,j)=k\}.$
The double indicator selects exactly one expert–advice pair per instance. The loss has the same structure as the standard deferral loss (1) but ranges over $J(K+1)$ actions rather than $J$, and it depends on the policy only through the executed pair (r(x),q(x,r(x)))(r(x),\,q(x,r(x))). When only the null advice action is available, the query function is trivial, $c_{j,0}(x,a,y)=c_{j}(x,y)$, and the loss reduces to $\ell_{\mathrm{def}}$ of (1); thus the standard deferral setting is a special case (Mozannar and Sontag, 2020; Mao et al., 2023a; Montreuil et al., 2026a).
4.3 Deferral with Advice Dominates Standard Deferral
We now derive the Bayes-optimal policy. Its structure reveals two properties: the choice of advice is inherently coupled to the choice of expert, and deferral with advice is more powerful than standard Learning-to-Defer.
The Bayes-optimal policy.
In Section 4.1 we argued that the value of advice is expert-dependent, and the policy of Definition 2 reflects this by conditioning the query $q(x,j)$ on the expert. We now show that minimizing the conditional risk of Definition 2 recovers the desired policy. To determine which expert is best, the system must first evaluate each expert at its best-advised cost—this requires solving the advice selection problem for every expert independently. The optimal query necessarily depends on the expert, since the best advice for expert $j$ is generally different from the best advice for expert $j^{\prime}$. The router then compares experts on their optimally-advised costs rather than their unaided ones.
Lemma 3 (Bayes-optimal deferral-advice policy).
For $\mathbb{P}_{X}$-a.e. $x$, the policy $(r^{\star},q^{\star})$ that minimizes $\mathbb{E}[\ell_{\mathrm{def\text{-}adv}}(r,q;\,X,A,Y,\mathbf{e})]$ over all measurable policies satisfies:
  1. 1.
    The Bayes query selects the best advice for each expert:
    $q^{\star}(x,j)\;\in\;\operatorname*{arg\,min}_{k\in[K]_{0}}\;\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr].$ (5)
  2. 2.
    The Bayes router selects the best optimally-advised expert:
    $r^{\star}(x)\;\in\;\operatorname*{arg\,min}_{j\in[J]}\;\min_{k\in[K]_{0}}\;\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr].$ (6)
The proof is given in Appendix B.1. The loss depends on the policy only through the executed pair (r(x),q(x,r(x)))(r(x),\,q(x,r(x))), so the population risk can be minimized pointwise in $x$ over the finite cost table. The result reveals a fundamental coupling: the router does not compare experts on their unaided costs $\mathbb{E}[c_{j}(X,Y)\mid X=x]$ as in the standard setting, but on their optimally-advised ones. An expert that would never be selected under standard deferral may become optimal once paired with the right advice. The two decisions—who acts and what they see—cannot be separated without loss.
When is advice worth acquiring?
In any practical system, acquiring information has a cost, and a sensible policy should query only when the expected benefit justifies the expense. The Bayes-optimal query function satisfies exactly this principle. Under the cost decomposition (4), the system acquires advice $k\geq 1$ for expert $j$ if and only if the expected reduction in task loss exceeds the advice fee.
Lemma 4 (Advice acquisition condition).
Under the cost decomposition (4), the Bayes-optimal query satisfies $q^{\star}(x,j)\neq 0$ if and only if there exists $k\in[K]$ such that
$\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(0)}),Y)\mid X=x\bigr]-\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(k)}),Y)\mid X=x\bigr]\;>\;\gamma_{k}.$ (7)
When no source satisfies this condition, $q^{\star}(x,j)=0$ and the expert acts on $x$ alone.
The proof is given in Appendix B.2. The left-hand side of (7) measures how much advice $k$ improves expert $j$’s prediction; the right-hand side is the price of obtaining it. The system acquires advice only when it is expected to pay for itself.
Deferral with advice dominates L2D.
What does this additional flexibility buy us at the population level? At a minimum, it cannot hurt. The no-advice action $k=0$ remains available, so the learner can always recover the standard L2D decision if acquiring advice is not worth its cost. The point, however, is stronger than a simple nesting observation. For each expert, the best advised action is never worse than the unaided one; once this pointwise improvement is propagated through the comparison over experts and the expectation over $X$, the Bayes risk of deferral with advice can only improve upon that of standard deferral.
Lemma 5 (Deferral with advice dominates standard deferral).
For every expert $j\in[J]$ and $\mathbb{P}_{X}$-a.e. $x$,
$\min_{k\in[K]_{0}}\;\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr]\;\leq\;\mathbb{E}\bigl[c_{j,0}(X,A,Y)\mid X=x\bigr].$ (8)
Consequently, if $\mathcal{E}^{\star}_{\ell_{\mathrm{def}}}$ denotes the Bayes risk of the standard deferral setting of Section 3, then
$\mathcal{E}^{\star}_{\ell_{\mathrm{def\text{-}adv}}}\;\leq\;\mathcal{E}^{\star}_{\ell_{\mathrm{def}}}.$ (9)
Equality holds if and only if, for $\mathbb{P}_{X}$-a.e. $x$,
$\min_{j\in[J]}\min_{k\in[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr]=\min_{j\in[J]}\mathbb{E}\bigl[c_{j,0}(X,A,Y)\mid X=x\bigr].$
The proof is given in Appendix B.3. By Lemma 4, the inequality is strict whenever advice lowers the best achievable conditional cost on a set of positive $\mathbb{P}_{X}$-measure. Thus advice extends L2D not only formally, but also in terms of achievable optimal risk.
4.4 The Separated Surrogate Fails
Because the protocol is sequential, the natural first attempt is to learn it sequentially as well: one router score, one query head per expert, and smooth logistic replacements for the hard indicators (Bartlett et al., 2006). But does mirroring the control flow suffice for consistency? It does not. We show that a broad family of separated surrogates already fails in the smallest non-trivial case: two experts $(J=2)$ and one binary advice $(K=1)$.
Let $s_{r}^{b}:\mathcal{X}\to\mathbb{R}$ be a binary router score and, for each expert $j\in\{1,2\}$, let $s_{q_{j}^{b}}:\mathcal{X}\to\mathbb{R}$ be a binary query score. Writing $\mathbf{s}_{q^{b}}=(s_{q_{1}^{b}},s_{q_{2}^{b}})$, the induced decisions are $r^{b}(x)=1+\mathbf{1}\{s_{r}^{b}(x)\geq 0\}$ and $q_{j}^{b}(x)=\mathbf{1}\{s_{q_{j}^{b}}(x)\geq 0\}$.
The natural separated construction.
In standard learning theory, one replaces margin indicators by smooth convex upper bounds (Zhang, 2002; Bartlett et al., 2006; Steinwart, 2007). Applying this idea to the binary executed-pair loss yields the logistic envelopes $\Phi_{0}(u)=\log(1+e^{-u})$ and $\Phi_{1}(u)=\log(1+e^{u})$. Rather than analyzing only the basic surrogate obtained by directly substituting these envelopes, we study a strictly larger family that encompasses every natural modification a practitioner might attempt: monotone cost transforms, reparameterized query scores, and flexible positive router weights. Let $\nu:[0,C]\to(0,\infty)$ be a monotone cost transform, $G:\mathbb{R}\to(0,\infty)$ a query amplitude, and $U:\mathbb{R}\to\mathbb{R}$ a strictly increasing query reparameterization. The resulting router-query separated surrogate family is
$\Phi_{\mathrm{def\text{-}adv}}^{r/q}(s_{r}^{b},\mathbf{s}_{q^{b}})\;=\;\sum_{j=1}^{2}\sum_{k=0}^{1}\Psi_{j}(s_{r}^{b}(x))\nu(c_{j,k}(x,a,y))\,G(s_{q_{j}^{b}}(x))\,\Phi_{k}(U(s_{q_{j}^{b}}(x))).$ (10)
Setting $\nu=\mathrm{id}$, $G\equiv 1$, $U=\mathrm{id}$, and $\Psi_{j}=\Phi_{j-1}$ recovers the basic construction obtained by replacing each indicator in the true loss with its logistic envelope—the standard recipe in cost-sensitive surrogate design (Cortes et al., 2016, 2024a); we give the full derivation in Appendix B.4.1. The broader family ensures that the negative result below is not due to specific parameterization.
Where the mismatch enters.
This separation has one decisive consequence. After profiling over the query heads, each expert is summarized by a scalar quantity
$F(u,v)\;\coloneqq\;\inf_{t\in\mathbb{R}}\;\bigl[\nu(u)\,G(t)\,\Phi_{0}(U(t))+\nu(v)\,G(t)\,\Phi_{1}(U(t))\bigr]$ (11)
that depends on the entire row, not only on its best entry. The Bayes rule, by contrast, compares row minima: $\min\{c_{1,0},c_{1,1}\}$ versus $\min\{c_{2,0},c_{2,1}\}$ (cf. (6)). The surrogate and the Bayes rule compress each row in fundamentally different ways. That difference is the entire obstruction.
We formalize this failure at the level of Fisher consistency, the weakest pointwise requirement that every surrogate minimizer decode to a Bayes-optimal action (Zhang, 2002; Bartlett et al., 2006; Steinwart, 2007; Tewari and Bartlett, 2007).
Theorem 6 (Broad Fisher inconsistency of separated router/query surrogates).
Under mild regularity conditions on $\nu$, $G$, $U$, $\Psi_{1}$, $\Psi_{2}$ (Appendix B.4), for every $b\in(0,C)$ and $\varepsilon\in(0,C-b)$, there exists $\delta\in(0,b)$ such that the pointwise cost table
$\mathbf{c}(x,a,y)\;=\;\begin{pmatrix}b-\delta&C\\ b&b+\varepsilon\end{pmatrix}$ (12)
has Bayes-optimal expert $j^{\star}=1$, while the unique minimizer of the pointwise surrogate (10) decodes to expert $2$. In particular, the entire family (10) is not Fisher consistent for the true protocol loss.
Expert $1$ has one excellent action and one catastrophic one, whereas expert $2$ has two mediocre actions. Bayes compares row minima, ignores the catastrophic entry, and therefore prefers expert $1$. The separated surrogate behaves differently. After profiling out the query head, row $1$ is summarized by $F(b-\delta,C)$, which still depends on the catastrophic entry. Appendix B.4.4 shows that $F(u,v)$ is strictly increasing in $v$ for fixed $u\in(0,C)$, so this distortion is unavoidable. The router therefore compares profiled row summaries rather than row minima and selects expert $2$.
Because the contradiction already appears in the smallest non-trivial setting, it cannot be dismissed as an artifact of a particular parameterization or of a large action space. Appendix C.2 complements this theorem with a theorem-aligned synthetic experiment, where the separated surrogate’s excess risk stays bounded away from zero and remains substantially worse than the augmented construction introduced next.
4.5 The Augmented Surrogate
The separated surrogate of Section 4.4 fails because profiling the query heads row by row distorts the router’s comparison. This raises the central question of whether the deferral-advice problem admits a calibrated surrogate at all, given that the separated construction is already inconsistent. We answer this question in the affirmative.
Composite actions.
Recall that the loss depends only on the executed pair (r(x),q(x,r(x)))(r(x),\,q(x,r(x))). Define the composite action set $\Pi\coloneqq[J]\times[K]_{0}$, of cardinality $|\Pi|=J(K+1)$. Each action $(j,k)\in\Pi$ represents the joint decision to route to expert $j$ and acquire advice $k$. A composite predictor is a measurable map $\pi:\mathcal{X}\to\Pi$. Let $\mathcal{H}_{\pi}\subseteq\{s_{\pi}:\mathcal{X}\times\Pi\to\mathbb{R}\}$ be a hypothesis class of measurable scoring functions; each $s_{\pi}\in\mathcal{H}_{\pi}$ induces a composite predictor via $\pi(x)=\operatorname*{arg\,max}_{(j,k)\in\Pi}s_{\pi}(x,(j,k))$. We write $\mathbf{s}_{\pi}(x)=(s_{\pi}(x,(j,k)))_{(j,k)\in\Pi}$ for the full score vector.
This reformulation does not alter the underlying decision problem. A sequential policy $(r,q)$ induces the composite action $\pi(x)=(r(x),\,q(x,r(x)))$, and every composite action $(j,k)$ can in turn be implemented by a sequential policy with the same realized loss. The composite view simply makes explicit what the loss has depended on all along: the executed expert–advice pair. Proposition 16, proved in Appendix B.5, shows that the sequential and composite formulations therefore induce the same population risks and share the same Bayes-optimal decisions as those identified in Lemma 3.
The augmented surrogate.
The composite view aligns the learning problem with the Bayes comparison itself. One still faces the non-differentiability of the true loss from Definition 2, so the remaining task is to construct a surrogate on the augmented action space. The key step, proved in Appendix B.6, is a decomposition of the realized loss into an action-independent offset and a weighted sum of composite mismatch indicators $\mathbf{1}\{\pi(x)\neq(j,k)\}$. Since the offset does not affect the pointwise comparison across actions, surrogate design can focus entirely on the mismatch term. The resulting problem is a weighted multiclass prediction task over $\Pi$.
More precisely, Appendix B.6 proves that
$\ell_{\mathrm{def\text{-}adv}}(\pi;\,x,a,y,\mathbf{e})\;=\;D(x,a,y)+\sum_{i\in\Pi}w_{i}(x,a,y)\,\mathbf{1}\{\pi(x)\neq i\},$ (13)
where $D(x,a,y)$ is independent of $\pi$ and $w_{i}(x,a,y)\;\coloneqq\;\max_{i^{\prime}\in\Pi}c_{i^{\prime}}(x,a,y)-c_{i}(x,a,y).$
To surrogate this mismatch term, we draw on the comp-sum family of Mao et al. (2023b). For $\tau\geq 0$, define the multiclass loss
$\Phi^{\tau}_{01}(\mathbf{s}_{\pi}(x),i)\;\coloneqq\;\begin{cases}\frac{1}{1-\tau}\left(\left[\sum_{i^{\prime}\in\Pi}e^{s_{\pi}(x,i^{\prime})-s_{\pi}(x,i)}\right]^{1-\tau}-1\right),&\tau\neq 1,\\ \log\left(\sum_{i^{\prime}\in\Pi}e^{s_{\pi}(x,i^{\prime})-s_{\pi}(x,i)}\right),&\tau=1.\end{cases}$ (14)
This parameterized family recovers sum-exponential (Weston and Watkins, 1998), logistic (Ohn Aldrich, 1997), generalized cross-entropy (Zhang and Sabuncu, 2018), and MAE (Ghosh et al., 2017). It therefore provides a unified way to instantiate the augmented surrogate.
Applying (14) to the weighted mismatch term yields the augmented deferral-advice surrogate.
Lemma 7 (Augmented deferral-advice surrogate).
For every scoring function $s_{\pi}$ and realization $(x,a,y)$, define the augmented deferral-advice surrogate by
$\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}(s_{\pi};\,x,a,y,\mathbf{e})\;\coloneqq\;\sum_{i\in\Pi}w_{i}(x,a,y)\;\Phi^{\tau}_{01}(\mathbf{s}_{\pi}(x),\,i),$ (15)
with weights $w_{i}(x,a,y)$ defined in (13).
Consistency: the main positive result.
Lemma 7 gives the right starting point, but it is not the whole story. Simply replacing the mismatch indicators by a smooth loss is not enough. The separated construction of Section 4.4 is precisely a counterexample: despite its natural form, Theorem 6 shows that its minimizers can still decode to the wrong executed decision. What matters is whether the surrogate preserves the comparison that the protocol actually makes.
We now establish the central theoretical guarantee of the paper: the augmented surrogate does not merely avoid inconsistency, it admits a quantitative excess-risk transfer bound for the entire comp-sum family and hence for every $\tau\geq 0$.
Theorem 8 ($\mathcal{H}_{\pi}$-consistency of the augmented surrogate).
For every $\tau\geq 0$, the augmented surrogate $\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}$ satisfies an $\mathcal{H}_{\pi}$-consistency bound with respect to $\ell_{\mathrm{def\text{-}adv}}$. Specifically, for every $s_{\pi}\in\mathcal{H}_{\pi}$ and its induced composite predictor $\pi$,
$\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\pi)-\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}^{B}(\mathcal{H}_{\pi})+\mathcal{U}_{\ell_{\mathrm{def\text{-}adv}}}(\mathcal{H}_{\pi})\leq\widetilde{\Gamma}_{\tau}\!\left(\mathcal{E}_{\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}}(s_{\pi})-\mathcal{E}_{\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}}^{\ast}(\mathcal{H}_{\pi})+\mathcal{U}_{\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}}(\mathcal{H}_{\pi})\right).$ (16)
Here, $\widetilde{\Gamma}_{\tau}(u)\;\coloneqq\;\mathbb{E}\bigl[\|\mathbf{w}(X)\|_{1}\bigr]\,\Gamma_{\tau}\!\left(\frac{u}{\mathbb{E}[\|\mathbf{w}(X)\|_{1}]}\right),$ where $\Gamma_{\tau}$ is the non-negative, non-decreasing, concave transfer function of the multiclass $\mathcal{H}_{\pi}$-consistency bound for $\Phi^{\tau}_{01}$ on $\Pi$, and $\mathbf{w}(x)\coloneqq(\mathbb{E}[w_{i}(X,A,Y)\mid X=x])_{i\in\Pi}$. For the logsoftmax surrogate $(\tau=1)$, we have $\Gamma_{1}(u)=\sqrt{2u}$ (Mao et al., 2023b).
We prove this theorem in Appendix B.8. When no advice is available ($K=0$, so $\Pi=[J]$), the bound (16) recovers the $\mathcal{H}_{r}$-consistency bounds established for standard Learning-to-Defer by Mao et al. (2023a, 2024c); Montreuil et al. (2025b, 2026a). The left-hand side measures the true excess risk up to minimizability gaps; the right-hand side is controlled by the surrogate excess risk through $\widetilde{\Gamma}_{\tau}$, which inherits the properties of the base comp-sum bound on $\Pi$.
Corollary 9 (Asymptotic Bayes-risk consistency).
If the minimizability gaps in Theorem 8 vanish and $\mathcal{E}_{\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}}(\widehat{s}_{\pi,n})-\mathcal{E}_{\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}}^{\ast}(\mathcal{H}_{\pi})\;\to\;0$ in probability, then the induced predictors $\widehat{\pi}_{n}$ satisfy $\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\widehat{\pi}_{n})-\inf_{\pi}\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\pi)\;\to\;0$ in probability.
Corollary 9 is the statistical payoff: whenever the hypothesis class is rich enough and the surrogate is well-optimized, the induced policy converges to the Bayes-optimal deferral-advice risk (Steinwart, 2007). When $\mathcal{H}_{\pi}$ contains all measurable functions, both minimizability gaps vanish by definition (Zhang and Agarwal, 2020; Bartlett et al., 2006; Long and Servedio, 2013), and the bound reduces to classical Bayes-consistency. The separated surrogate offers no such guarantee—its minimizers can converge to a provably suboptimal policy regardless of model capacity.
5 Experiments
We evaluate the augmented surrogate on the deployment objective of the paper, namely the true deferral-advice loss. In the main paper we focus on FEVER, where expert-dependent advice arises naturally because heterogeneous models benefit differently from retrieved evidence. The appendix gives the full FEVER protocol together with additional tabular, multi-modal, and synthetic experiments; see Appendix C.3, Appendix C.4, Appendix C.5, and Appendix C.2.
5.1 FEVER setup
Task, experts, and advice.
We use the FEVER fact-verification benchmark (Thorne et al., 2018). The expert pool contains Qwen3-4B-Instruct (Yang et al., 2025), Qwen3-8B (Yang et al., 2025), and DeBERTa-v3-large (He et al., 2021), with consultation costs $\beta=(0.03,0.05,0.04)$. Advice consists of retrieved evidence. The five actions are no retrieval, BM25 top-$1$, top-33, top-55, and a query-reformulation pipeline produced by Qwen2.5-1.5B. The base advice costs are $\gamma^{\mathrm{base}}=(0,0.015,0.02,0.03,0.01)$, and the deployed cost is $\gamma_{k}=\lambda\gamma_{k}^{\mathrm{base}}$ for $\lambda\in\{0,0.5,1,2,4,6,8,10\}$.
Training protocol and baselines.
For every training example and every executed pair $(j,k)$, we precompute the expert prediction and the realized cost $c_{j,k}(x,a,y)=\mathbf{1}\{e_{j}(x,\widetilde{a}^{(k)})\neq y\}+\beta_{j}+\gamma_{k}$. The composite scorer is an MLP trained with Adam; full architecture and hyperparameter details are in Appendix C.3. We evaluate the empirical validation average of $\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\pi)$. We compare against: (i) standard L2D (Mao et al., 2023a), restricted to $k=0$, so it can choose the expert but never acquire advice; (ii) the best fixed expert under its best advice; (iii) partial-randomization ablations: learned expert with random advice, and learned advice with random expert; (iv) joint-random policy over expert-advice pairs.
5.2 Results
Table 1: Validation accuracy (%) of each expert under each advice action at $\lambda=0$. Bold marks the best hint for each expert. The preferred hint is expert-dependent.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3 k=4k{=}4
(none) (top-1) (top-3) (top-5) (reform.)
Qwen3-4B-Instruct 33.5 33.4 32.9 36.3 34.0
Qwen3-8B 33.5 33.4 34.1 41.1 35.7
DeBERTa-v3-large 33.2 42.9 42.5 42.3 39.2
Table 1 shows that the value of advice is strongly expert-dependent. Without retrieval, all three experts operate near chance. As soon as evidence is provided, their preferred actions diverge: DeBERTa is best with top-$1$ retrieval, whereas both LLMs prefer top-$5$, and the larger LLM benefits substantially more from it. This is precisely the regime targeted by our theory.
Table 2: Validation true deferral-advice loss across advice-cost multipliers $\lambda$. Entries are mean $\pm$ standard deviation over random seeds. Lower is better. Bold marks the best mean in each row.
$\lambda$ Ours L2D Best fixed Learned $j$, random $k$ Random $j$, learned $k$ Random $(j,k)$
0 0.555 $\pm$ 0.002 0.619 $\pm$ 0.001 0.611 0.633 $\pm$ 0.005 0.640 $\pm$ 0.007 0.675 $\pm$ 0.006
$0.5$ 0.558 $\pm$ 0.002 0.619 $\pm$ 0.001 0.619 0.640 $\pm$ 0.005 0.657 $\pm$ 0.007 0.682 $\pm$ 0.006
11 0.566 $\pm$ 0.001 0.619 $\pm$ 0.001 0.626 0.647 $\pm$ 0.007 0.664 $\pm$ 0.007 0.690 $\pm$ 0.006
22 0.580 $\pm$ 0.001 0.619 $\pm$ 0.001 0.641 0.662 $\pm$ 0.006 0.674 $\pm$ 0.009 0.705 $\pm$ 0.006
44 0.596 $\pm$ 0.003 0.619 $\pm$ 0.001 0.671 0.695 $\pm$ 0.006 0.695 $\pm$ 0.007 0.735 $\pm$ 0.007
66 0.607 $\pm$ 0.002 0.619 $\pm$ 0.001 0.695 0.723 $\pm$ 0.006 0.704 $\pm$ 0.007 0.765 $\pm$ 0.007
88 0.613 $\pm$ 0.001 0.619 $\pm$ 0.001 0.695 0.760 $\pm$ 0.006 0.711 $\pm$ 0.006 0.795 $\pm$ 0.007
$10$ 0.619 $\pm$ 0.001 0.619 $\pm$ 0.001 0.695 0.788 $\pm$ 0.006 0.715 $\pm$ 0.008 0.825 $\pm$ 0.007
Table 2 gives the main conclusion. Our method achieves the lowest mean true loss for seven of the eight cost regimes. The gain is largest when advice is cheap: at $\lambda=0$, it improves over L2D by $0.064$ in mean true risk. The advantage then shrinks smoothly as advice becomes more expensive and the optimal policy moves toward no-query decisions. The same transition appears in the best fixed baseline, which switches from DeBERTa with top-$1$ retrieval for $\lambda\leq 4$ to the cheapest no-query expert for $\lambda\geq 6$.
Two further points are worth emphasizing. First, our method remains below the best fixed pair throughout the sweep, showing that the benefit is not merely access to more actions but input-dependent selection of the executed pair. Second, both partial-randomization ablations are substantially worse than the full policy. Their relative ordering also changes with the cost regime: at low advice cost, randomizing the expert is slightly more damaging, whereas at high advice cost, randomizing the advice becomes worse. This shows that the gain does not come from only one half of the decision. Moreover, the baseline with random expert and learned advice consistently improves over the fully random policy, indicating that the learned advice component is not a global preference for one retrieval action but a meaningful expert-conditional decision. When advice becomes too expensive ($\lambda=10$), our method ties the L2D baseline.
6 Conclusion
We introduced Learning-to-Defer with advice, a setting in which the system must decide not only which expert should act, but also which additional information that expert should receive. We show that a broad family of natural separated router-query surrogates can be Fisher inconsistent even in the smallest non-trivial setting. We gave an augmented surrogate that learns directly over the composite expert-advice action space and proved an $\mathcal{H}$-consistency guarantee. Empirically, the resulting method improves over standard Learning-to-Defer across language, tabular, and multi-modal settings, while adapting its advice-acquisition behavior to the cost regime. More broadly, these results show that once information acquisition is expert dependent, Learning-to-Defer must reason not only about who acts, but also about what they get to see.
Acknowledgment
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2023-01-041-J) and by A*STAR, and is part of the programme DesCartes which is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and Section 1 state four contributions (formulation, Fisher inconsistency, augmented surrogate with $\mathcal{H}$-consistency, experiments). Each is supported by formal results (Lemma 1, Theorem 1, Theorem 2, Corollary 1) and experiments (Section 5, Appendices B–E).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The Conclusion contains a dedicated Limitations paragraph discussing the full-information training assumption, the growth of the composite action space, and the population-level nature of the consistency guarantees.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate “Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All lemmas, theorems, and corollaries are numbered and cross-referenced. Full proofs with explicit regularity conditions are given in Appendix B. The main text provides proof intuition for every result.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 5 describes the task, experts, advice actions, cost structure, and baselines. Appendix C describes the global experimental protocol including model architecture, optimizer, and hyperparameters. Appendices D–E give per-experiment details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Code will be released upon acceptance. The FEVER dataset is publicly available. All experimental details needed for reimplementation are provided in the appendix.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 5 summarizes the main experimental setup. Full details including architecture, optimizer, learning rate, number of seeds, and data splits are given in Appendix C (global protocol) and Appendices D–E (per-experiment).
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Table 2 reports mean $\pm$ standard deviation over multiple random seeds for all learned methods. The caption states that entries are mean $\pm$ standard deviation over random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer “Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix C describes the compute infrastructure used for experiments, including hardware type and approximate runtimes.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research uses publicly available datasets and open-source models. No human subjects are involved. The work conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This work is foundational research on surrogate loss design for Learning-to-Defer. It does not target a specific application domain and has no direct path to negative societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not release pretrained models, datasets, or other assets that pose a risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and models used are cited. FEVER is released under a Creative Commons license. The Qwen and DeBERTa models are open-source and properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new datasets or models.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve human subjects research.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLMs (Qwen3-4B-Instruct, Qwen3-8B) are used as experts in the FEVER experiment and Qwen2.5-1.5B is used for query reformulation. Their usage is described in Section 5 and Appendix D.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
Contents
Appendix A Algorithms and Visualization
A.1 Notation Summary
Table 3 collects the main symbols used throughout the paper. We keep only the objects that are central to the formulation and theory. As in the main text, $(j,k)$ denotes an executed expert–advice pair, while $i\in\Pi$ is used as shorthand for a composite action when convenient.
Table 3: Notation summary.
Symbol Meaning
Data, experts, and advice
$(X,Y)\sim\mathcal{D}$ Input-output random pair.
$(x,y)$ One realization of $(X,Y)$.
$A=(A^{1},\dots,A^{K})$ Advice random vector.
$(x,a,y)$ One realization of the full deferral-advice instance.
$J]=\{1,\dots,J\}$ Expert index set.
$K]_{0}=\{0,1,\dots,K\}$ Advice-action set, where $k=0$ denotes no advice.
$m_{k}$ Masking operator that reveals at most advice source $k$.
$\widetilde{a}^{(k)}$ Masked realized advice under action $k$.
$\widetilde{A}^{(k)}$ Masked advice random variable under action $k$.
$e_{j}$ Expert predictor for expert $j$.
$\mathbf{e}=(e_{1},\dots,e_{J})$ Tuple of expert predictors.
$\psi$ Task loss.
$\beta_{j}$ Expert consultation cost for expert $j$.
$\gamma_{k}$ Advice acquisition cost for advice action $k$.
$c_{j}(x,y)$ Standard L2D cost of routing $(x,y)$ to expert $j$.
$c_{j,k}(x,a,y)$ Deferral-advice cost of executing expert–advice pair $(j,k)$.
Policies, actions, and risks
$r:\mathcal{X}\to[J]$ Router in standard Learning-to-Defer.
$q:\mathcal{X}\times[J]\to[K]_{0}$ Expert-conditional query function.
$\ell_{\mathrm{def}}$ Standard deferral loss.
$\ell_{\mathrm{def\text{-}adv}}$ True deferral-advice loss.
$\mathcal{E}_{\ell_{\mathrm{def}}}$ Population risk associated with $\ell_{\mathrm{def}}$.
$\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}$ Population risk associated with $\ell_{\mathrm{def\text{-}adv}}$.
$r^{\star}$ Bayes-optimal router.
$q^{\star}$ Bayes-optimal query function.
$\Pi=[J]\times[K]_{0}$ Composite action space of executed expert–advice pairs.
$i\in\Pi$ Composite action index, typically identified with a pair $(j,k)$.
$\pi:\mathcal{X}\to\Pi$ Composite predictor.
Surrogates and consistency quantities
$\mathcal{H}_{r}$ Hypothesis class for router scores.
$\mathcal{H}_{q}$ Hypothesis class for query scores.
$\mathcal{H}_{\pi}$ Hypothesis class for composite scores.
$s_{r}$ Router score function.
$\mathbf{s}_{r}(x)$ Router score vector over experts.
$s_{q}$ Query score function over expert–advice pairs $(j,k)$.
$\mathbf{S}_{q}(x)$ Query-score matrix with entries $s_{q}(x,j,k)$.
$s_{\pi}$ Composite score function.
$\mathbf{s}_{\pi}(x)$ Composite score vector over $\Pi$.
$\Phi_{\mathrm{def\text{-}adv}}^{r/q}$ Separated router/query surrogate family.
$F(u,v)$ Profiled row summary in the separated-surrogate analysis.
$D(x,a,y)$ Action-independent offset in the composite-loss decomposition.
$w_{i}(x,a,y)$ Mismatch weight associated with composite action $i$.
$\Phi^{\tau}_{01}$ Comp-sum multiclass loss indexed by $\tau\geq 0$.
$\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}$ Augmented deferral-advice surrogate.
$\Gamma_{\tau}$ Transfer function for the base comp-sum loss on $\Pi$.
$\widetilde{\Gamma}_{\tau}$ Transfer function in the augmented-surrogate consistency bound.
A.2 Visualization
Input $x$ Router scores $\mathbf{s}_{r}(x)\in\mathbb{R}^{J}$ Query-score matrix $\mathbf{S}_{q}(x)=\bigl(s_{q}(x,j,k)\bigr)_{j,k}$ Routed expert $r(x)=\operatorname*{arg\,max}_{j\in[J]}s_{r}(x,j)$ Advice index per expert rowwise argmax of $\mathbf{S}_{q}(x)$ $\bigl(q(x,1),\ldots,q(x,J)\bigr)$ Executed pair $\bigl(r(x),\,q(x,r(x))\bigr)$ only this pair enters the loss Revealed advice $\widetilde{a}^{(q(x,r(x)))}$ Executed decision expert $r(x)$ uses $\widetilde{a}^{(q(x,r(x)))}$ Output Prediction / decision
Figure 1: The deferral-advice protocol. From the input $x$, the router produces expert scores and the query model produces a $J\times(K+1)$ query-score matrix. Taking the argmax along each row yields one advice index per expert, but only the routed row is ever executed: the protocol uses the pair $\bigl(r(x),\,q(x,r(x))\bigr)$, reveals the corresponding masked advice, and executes expert $r(x)$ under that advice. When $q(x,r(x))=0$, the revealed advice is the null masked advice $\widetilde{a}^{(0)}$.
A.3 Algorithms
Algorithms 1, 3, and 2 summarize the two decision rules discussed in the paper. The first two are the proposed augmented approach: training from fully observed executed-pair costs, and deployment-time execution of the learned composite policy. Algorithm 2 then records the inference rule of the separated router/query parameterization, which is the natural sequential baseline analyzed in Section 4.4.
Algorithm 1 Training with the augmented surrogate
0: Training set $\{(x_{n},a_{n},y_{n})\}_{n=1}^{N}$, composite action set $\Pi=[J]\times[K]_{0}$, score model $s_{\pi}(\cdot,\cdot;\theta)$, comp-sum parameter $\tau\geq 0$, learning rate $\eta$
1: Initialize parameters $\theta$
2:repeat
3:  Sample a minibatch $B\subseteq\{1,\dots,N\}$
4:  for all $n\in B$ do
5:   for all $i=(j,k)\in\Pi$ do
6:    Compute the realized cost $c_{i}(x_{n},a_{n},y_{n})\equiv c_{j,k}(x_{n},a_{n},y_{n})$
7:   end for
8:   Set $w_{i}(x_{n},a_{n},y_{n})\leftarrow\max_{i^{\prime}\in\Pi}c_{i^{\prime}}(x_{n},a_{n},y_{n})-c_{i}(x_{n},a_{n},y_{n})$ for all $i\in\Pi$
9:   Form the sample loss
$L_{n}(\theta)\leftarrow\sum_{i\in\Pi}w_{i}(x_{n},a_{n},y_{n})\,\Phi^{\tau}_{01}(\mathbf{s}_{\pi}(x_{n};\theta),i)$
10:  end for
11:  Update $\theta\leftarrow\theta-\eta\nabla_{\theta}\left(|B|^{-1}\sum_{n\in B}L_{n}(\theta)\right)$
12:until convergence
13:return learned parameters $\widehat{\theta}$
Algorithm 2 Inference with a separated router/query policy
0: Input $x$, router score model $s_{r}(\cdot,\cdot)$, query score model $s_{q}(\cdot,\cdot,\cdot)$
1: Form the query-score matrix
$\mathbf{S}_{q}(x)\leftarrow\bigl(s_{q}(x,j,k)\bigr)_{j\in[J],\,k\in[K]_{0}}\in\mathbb{R}^{J\times(K+1)}$
2: For each expert $j$, take the argmax along row $j$ of $\mathbf{S}_{q}(x)$:
$\widehat{k}_{j}(x)\in\operatorname*{arg\,max}_{k\in[K]_{0}}s_{q}(x,j,k)$
3: The routed expert uses its own rowwise maximizer
4: Compute the routed expert
$\widehat{j}(x)\in\operatorname*{arg\,max}_{j\in[J]}s_{r}(x,j)$
5: Select the advice actually executed by the routed expert
$\widehat{k}(x)\leftarrow\widehat{k}_{\widehat{j}(x)}(x)$
6:if $\widehat{k}(x)=0$ then
7:  Set $\widetilde{a}^{(\widehat{k}(x))}\leftarrow\widetilde{a}^{(0)}$
8:else
9:  Acquire advice source $\widehat{k}(x)$ and reveal $\widetilde{a}^{(\widehat{k}(x))}$
10:end if
11: Route $x$ to expert $\widehat{j}(x)$ and output $e_{\widehat{j}(x)}(x,\widetilde{a}^{(\widehat{k}(x))})$
12:return routed expert $\widehat{j}(x)$, executed advice $\widehat{k}(x)$, and prediction $e_{\widehat{j}(x)}(x,\widetilde{a}^{(\widehat{k}(x))})$
Algorithm 3 Inference with the learned composite policy
[…]
1: Select the composite action
[…]
2:if $\widehat{k}=0$ then
3:  Set $\widetilde{a}^{(\widehat{k})}\leftarrow\widetilde{a}^{(0)}$
4:else
5:  Acquire advice source $\widehat{k}$ and reveal the masked advice $\widetilde{a}^{(\widehat{k})}$
6:end if
[…]
Algorithm 4 FEVER inference example
[…]
1: Define the FEVER expert set $\{\text{Qwen3-4B-Instruct},\text{Qwen3-8B},\text{DeBERTa-v3}\}$
2: Define the advice actions $\{0,\text{top-1},\text{top-3},\text{top-5},\text{reformulate+top-5}\}$
3: Select the executed pair
[…]
4:if $\widehat{k}=0$ then
5:  Set $\widetilde{a}^{(\widehat{k})}\leftarrow\widetilde{a}^{(0)}$ and issue no retrieval request
6:else if $\widehat{k}=\text{top-1}$ then
7:  Run BM25 on claim $x$ and keep the top-$1$ retrieved passage
8:  Mask all non-selected advice slots to form $\widetilde{a}^{(\widehat{k})}$
9:else if $\widehat{k}=\text{top-3}$ then
10:  Run BM25 on claim $x$ and keep the top-$3$ retrieved passages
11:  Mask all non-selected advice slots to form $\widetilde{a}^{(\widehat{k})}$
12:else if $\widehat{k}=\text{top-5}$ then
13:  Run BM25 on claim $x$ and keep the top-$5$ retrieved passages
14:  Mask all non-selected advice slots to form $\widetilde{a}^{(\widehat{k})}$
15:else
16:  Use the reformulation model to rewrite the claim query
17:  Mask all non-selected advice slots to form $\widetilde{a}^{(\widehat{k})}$
18:end if
19: Route the claim and masked advice to expert $\widehat{j}$
20: Output the FEVER label $e_{\widehat{j}}(x,\widetilde{a}^{(\widehat{k})})\in\{\textsc{Supports},\textsc{Refutes},\textsc{Not Enough Info}\}$
[…]
Appendix B Proofs
Example 1 (Running conditional-cost table).
Before turning to the formal derivations, it is useful to keep one concrete conditional cost table in mind. Fix an input $x$ and consider
$\bar{\mathbf{C}}(x)\;\coloneqq\;\bigl(\mathbb{E}[c_{j,k}(X,A,Y)\mid X=x]\bigr)_{j\in[3],\,k\in\{0,1,2\}}=\begin{pmatrix}0.32&0.41&0.36\\ 0.35&0.27&0.40\\ 0.38&0.33&0.24\end{pmatrix}$.
Each row corresponds to an expert, each column to an advice action, and $k=0$ denotes the no-advice option. The entries already include the entire executed cost: prediction error, expert fee, and advice fee. The query decision therefore acts row by row, because once the expert is fixed the learner still has to decide which advice source, if any, should be revealed to that expert. In this table, expert $1$ is best with no advice, expert $2$ is best with advice $1$, and expert $3$ is best with advice $2$. The expert-wise Bayes query therefore keeps only the row minima
(0.32, 0.27, 0.24),(0.32,\;0.27,\;0.24),
and the Bayes router then selects expert $3$, because $0.24$ is the smallest of these three values.
Standard Learning-to-Defer would reason differently. It would only compare the unaided column
(0.32, 0.35, 0.38),(0.32,\;0.35,\;0.38),
and would therefore select expert $1$. The richer protocol changes the decision because it allows each expert to be evaluated under its own best advice.
This example will be useful in the next three proofs. It illustrates the three main ideas: the best advice is expert-dependent, the router should compare experts only after this expert-wise query choice has been made, and the additional advice menu can strictly improve the achievable Bayes risk.
B.1 Proof of Lemma 3
See 3
Proof The key point is that the protocol pays only for the executed pair $(r(x),q(x,r(x)))$. Once this is made explicit, the Bayes problem reduces to a pointwise comparison of the conditional expected costs of the finitely many expert–advice pairs.
Indeed, by Definition 2, the population risk of any policy $(r,q)$ can be written as
$\mathbb{E}\bigl[\ell_{\mathrm{def\text{-}adv}}(r,q;\,X,A,Y,\mathbf{e})\bigr]=\mathbb{E}\Biggl[\sum_{j\in[J]}\sum_{k\in[K]_{0}}c_{j,k}(X,A,Y)\,\mathbf{1}\{r(X)=j\}\,\mathbf{1}\{q(X,j)=k\}\Biggr].$
Since exactly one pair $(j,k)$ is selected for each $X$, this simplifies to
$\mathbb{E}\bigl[c_{r(X),\,q(X,r(X))}(X,A,Y)\bigr].$
Thus the loss depends on the policy only through the executed pair. Applying the tower property of conditional expectation gives
$\mathbb{E}\bigl[c_{r(X),\,q(X,r(X))}(X,A,Y)\bigr]=\mathbb{E}\Bigl[\mathbb{E}\bigl[c_{r(X),\,q(X,r(X))}(X,A,Y)\mid X\bigr]\Bigr].$
Since $r(X)$ and $q(X,r(X))$ are deterministic functions of $X$, conditioning on $X$ fixes the pair (j,k)=(r(X),q(X,r(X)))(j,k)=(r(X),\,q(X,r(X))), and the inner expectation becomes $\mathbb{E}[c_{j,k}(X,A,Y)\mid X]$. It is therefore enough to minimize this conditional expectation pointwise in $x$.
For $\mathbb{P}_{X}$-a.e. $x$, the optimization problem is
$\min_{(j,k)\in[J]\times[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr].$
Since $J]\times[K]_{0}$ is a finite set and each $c_{j,k}$ is bounded (hence integrable), the conditional expectations $\mathbb{E}[c_{j,k}(X,A,Y)\mid X=x]$ are well-defined measurable functions of $x$. The minimum of finitely many measurable functions is measurable, so a measurable minimizer exists. This minimum can be decomposed as
$\min_{(j,k)\in[J]\times[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr]=\min_{j\in[J]}\;\min_{k\in[K]_{0}}\;\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr],$
so the optimal decision decomposes in the same way. For each expert $j$, the best query is any minimizer of the inner problem,
$q^{\star}(x,j)\in\operatorname*{arg\,min}_{k\in[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr].$
Once these expert-wise best query values are fixed, the router compares the resulting optimally-advised costs and therefore satisfies
$r^{\star}(x)\in\operatorname*{arg\,min}_{j\in[J]}\mathbb{E}\bigl[c_{j,q^{\star}(x,j)}(X,A,Y)\mid X=x\bigr]=\operatorname*{arg\,min}_{j\in[J]}\min_{k\in[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr].$
This is exactly the pair of Bayes decisions stated in (5) and (6). Example 2 shows this decomposition concretely: the query first selects one minimizing column per row, and the router then compares only these row minima. Consequently, the Bayes risk is
$\mathcal{E}^{\star}_{\ell_{\mathrm{def\text{-}adv}}}=\mathbb{E}\Bigl[\min_{j\in[J]}\;\min_{k\in[K]_{0}}\;\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X\bigr]\Bigr].$
Example 2 (Bayes query and Bayes router on a conditional cost table).
Fix an input $x$ and suppose the conditional executed-pair costs are
$\bar{\mathbf{C}}(x)\;\coloneqq\;\bigl(\mathbb{E}[c_{j,k}(X,A,Y)\mid X=x]\bigr)_{j\in[3],\,k\in\{0,1,2\}}=\begin{pmatrix}0.32&0.41&0.36\\ 0.35&0.27&0.40\\ 0.38&0.33&0.24\end{pmatrix}$.
Each row corresponds to one expert, and each column to one advice action, with $k=0$ denoting no advice. The Bayes query acts row by row: for expert $1$, the smallest entry is $0.32$, so $q^{\star}(x,1)=0$; for expert $2$, the smallest entry is $0.27$, so $q^{\star}(x,2)=1$; and for expert $3$, the smallest entry is $0.24$, so $q^{\star}(x,3)=2$. Once these query decisions are fixed, the router does not compare the full rows anymore. It compares the best achievable cost of each expert,
$\bigl(\mathbb{E}[c_{1,q^{\star}(x,1)}\mid X=x],\,\mathbb{E}[c_{2,q^{\star}(x,2)}\mid X=x],\,\mathbb{E}[c_{3,q^{\star}(x,3)}\mid X=x]\bigr)=(0.32,\;0.27,\;0.24),$
and therefore selects
$r^{\star}(x)=3.$
This is the content of Lemma 3: first determine, for each expert, which advice makes that expert cheapest; then route to the expert whose best-advised conditional cost is smallest.
B.2 Proof of Lemma 4
See 4
Proof The lemma formalizes the economic meaning of the Bayes query. Once the expert is fixed, the only remaining decision is whether the gain from revealing advice justifies its acquisition cost. The proof makes this trade-off explicit.
Fix an expert $j\in[J]$ and an input $x$. For every $k\in[K]_{0}$,
$\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr]=\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(k)}),Y)\mid X=x\bigr]+\beta_{j}+\gamma_{k}.$
The expert fee $\beta_{j}$ does not depend on the advice index and therefore plays no role in the comparison across $k$. For a fixed expert, the Bayes query thus solves
$q^{\star}(x,j)\;\in\;\operatorname*{arg\,min}_{k\in[K]_{0}}\;\Bigl\{\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(k)}),Y)\mid X=x\bigr]+\gamma_{k}\Bigr\}.$
The no-advice action corresponds to $k=0$ and has cost
$\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(0)}),Y)\mid X=x\bigr],$
because $\gamma_{0}=0$. The policy therefore prefers a queried advice source $k\geq 1$ to no advice exactly when
$\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(k)}),Y)\mid X=x\bigr]+\gamma_{k}\;<\;\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(0)}),Y)\mid X=x\bigr],$
which rearranges to
$\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(0)}),Y)\mid X=x\bigr]-\mathbb{E}\bigl[\psi(e_{j}(X,\widetilde{A}^{(k)}),Y)\mid X=x\bigr]\;>\;\gamma_{k}.$
This is exactly condition (7). Hence $q^{\star}(x,j)\neq 0$ if and only if at least one advice source yields an expected reduction in task loss that strictly exceeds its fee. If no such source exists, then every queried action is more expensive than its expected benefit, and the Bayes-optimal decision is $q^{\star}(x,j)=0$. This is precisely the claim of the lemma.
B.3 Proof of Lemma 5
See 5
Proof The intuition is simple. Deferral with advice enlarges the menu of admissible actions, but it never removes the standard L2D action. For any fixed expert, the best cost achievable with advice can therefore only improve upon, or match, the unaided one. The content of the lemma is that this row-wise comparison is exactly the one the Bayes router uses, so it lifts directly to the population Bayes risk.
Fix an expert $j\in[J]$ and an input $x$ in a full-measure set on which the relevant conditional expectations are defined. Since the null action $k=0$ belongs to $K]_{0}$, minimizing over the larger advice set immediately gives
$\min_{k\in[K]_{0}}\;\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr]\;\leq\;\mathbb{E}\bigl[c_{j,0}(X,A,Y)\mid X=x\bigr].$ (17)
This is exactly the pointwise statement (8): expert $j$ can always recover its unaided behavior by choosing $k=0$.
Because (17) holds for every expert, it remains true after taking the outer minimum over $j\in[J]$:
$\min_{j\in[J]}\min_{k\in[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr]\;\leq\;\min_{j\in[J]}\mathbb{E}\bigl[c_{j,0}(X,A,Y)\mid X=x\bigr].$
The left-hand side is precisely the conditional quantity minimized by the Bayes router of Lemma 3. On the right-hand side, the action $k=0$ coincides with standard Learning-to-Defer, so $c_{j,0}(x,a,y)=c_{j}(x,y)$ and therefore
$\mathbb{E}\bigl[c_{j,0}(X,A,Y)\mid X=x\bigr]=\mathbb{E}\bigl[c_{j}(X,Y)\mid X=x\bigr].$
Integrating over $X$ therefore yields
$\mathcal{E}^{\star}_{\ell_{\mathrm{def\text{-}adv}}}=\mathbb{E}\Bigl[\min_{j\in[J]}\min_{k\in[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X\bigr]\Bigr]\;\leq\;\mathbb{E}\Bigl[\min_{j\in[J]}\mathbb{E}\bigl[c_{j}(X,Y)\mid X\bigr]\Bigr]=\mathcal{E}^{\star}_{\ell_{\mathrm{def}}},$
which proves (9).
The equality case must be stated at the level of the best available expert, not expert by expert. Define
$g_{\mathrm{adv}}(x)\coloneqq\min_{j\in[J]}\min_{k\in[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr],\qquad g_{\mathrm{def}}(x)\coloneqq\min_{j\in[J]}\mathbb{E}\bigl[c_{j,0}(X,A,Y)\mid X=x\bigr].$
The first part of the proof shows that $g_{\mathrm{adv}}(x)\leq g_{\mathrm{def}}(x)$ for $\mathbb{P}_{X}$-a.e. $x$. Equivalently, the gap
$h(x)\coloneqq g_{\mathrm{def}}(x)-g_{\mathrm{adv}}(x)$
is nonnegative for $\mathbb{P}_{X}$-a.e. $x$. Equality in (9),
$\mathbb{E}[g_{\mathrm{adv}}(X)]=\mathbb{E}[g_{\mathrm{def}}(X)],$
is therefore equivalent to $\mathbb{E}[h(X)]=0$. Since $h(X)\geq 0$ almost surely, this happens if and only if
$g_{\mathrm{adv}}(x)=g_{\mathrm{def}}(x)\qquad\text{for }\mathbb{P}_{X}\text{-a.e.\ }x.$
This is exactly the condition stated in the lemma. Advice may improve some non-selected experts without changing the Bayes risk. What matters is whether it lowers the smallest conditional cost available at $x$ on a set of positive $\mathbb{P}_{X}$-measure.
Example 3 (Lemma 5 on the running table).
Return to Example 1. The conditional cost table is
$\bar{\mathbf{C}}(x)=\begin{pmatrix}0.32&0.41&0.36\\ 0.35&0.27&0.40\\ 0.38&0.33&0.24\end{pmatrix}$.
The Bayes query function is obtained row by row:
$q^{\star}(x,1)=0,\qquad q^{\star}(x,2)=1,\qquad q^{\star}(x,3)=2.$
This means that expert $1$ should act without advice, expert $2$ should receive advice source $1$, and expert $3$ should receive advice source $2$. The query decision therefore acts within each row: once the expert is fixed, it keeps only the advice action that makes that expert cheapest. After this expert-wise selection, the three rows reduce to the best-achievable cost vector
$\bigl(\mathbb{E}[c_{1,q^{\star}(x,1)}\mid X=x],\,\mathbb{E}[c_{2,q^{\star}(x,2)}\mid X=x],\,\mathbb{E}[c_{3,q^{\star}(x,3)}\mid X=x]\bigr)=(0.32,\;0.27,\;0.24).$
The Bayes router therefore chooses
$r^{\star}(x)=3,$
because expert $3$ now has the smallest achievable conditional cost.
For comparison, standard Learning-to-Defer only sees the no-advice column,
$\bigl(\mathbb{E}[c_{1,0}\mid X=x],\,\mathbb{E}[c_{2,0}\mid X=x],\,\mathbb{E}[c_{3,0}\mid X=x]\bigr)=(0.32,\;0.35,\;0.38),$
so its Bayes router would select expert $1$. Denoting this standard router by $r^{\star}_{\mathrm{def}}$, we therefore have
$r^{\star}_{\mathrm{def}}(x)=1\qquad\text{whereas}\qquad r^{\star}(x)=3.$
The reason is exactly the one formalized in Lemma 5: once the advice menu is available, each row can only stay the same or improve, because $k=0$ remains feasible. In this example, the no-advice comparison vector
(0.32, 0.35, 0.38)(0.32,\;0.35,\;0.38)
is replaced by the best-advised vector
(0.32, 0.27, 0.24).(0.32,\;0.27,\;0.24).
This new vector is coordinatewise no larger, and it is strictly smaller in the second and third coordinates. What matters for the lemma is the smallest entry of these two vectors, because that entry determines the Bayes decision at $x$. Here the best achievable conditional cost drops from
$g_{\mathrm{def}}(x)=0.32\qquad\text{to}\qquad g_{\mathrm{adv}}(x)=0.24.$
This is why the richer protocol changes the selected expert and strictly lowers the Bayes-risk contribution at this input. Had advice improved only a non-selected row while leaving the minimum at $0.32$, the overall Bayes risk would have remained unchanged. This is precisely the distinction captured by the equality condition in Lemma 5.
B.4 Proof of Theorem 6
The proof has a simple architecture. We begin from the exact binary deferral-advice loss and show how the usual logistic upper-bound recipe leads naturally to a separated router/query surrogate. We then identify the quantity that this surrogate actually compares after the query heads are optimized out: each expert row is compressed into a profiled summary $F(c_{j,0},c_{j,1})$, not into its row minimum. The final step is to construct a bounded cost table on which these two comparisons disagree. This shows that the failure is not tied to one peculiar parameterization, but to the separated architecture itself.
We work in the binary setting $J=2$, $K=1$ introduced in Section 4.4. The surrogate (10) is parameterized by a cost transform $\nu$, query functions $G$ and $U$, the logistic pair $\Phi_{0}$, $\Phi_{1}$, and router weights $\Psi_{1}$, $\Psi_{2}$.
Binary scores and decoded decisions.
Let $s_{r}^{b}:\mathcal{X}\to\mathbb{R}$ be a binary router score and, for each expert $j\in\{1,2\}$, let $s_{q_{j}^{b}}:\mathcal{X}\to\mathbb{R}$ be a binary query score. Write $\mathbf{s}_{q^{b}}=(s_{q_{1}^{b}},s_{q_{2}^{b}})$. As in the main text, the decoded decisions are
$r^{b}(x)=1+\mathbf{1}\{s_{r}^{b}(x)\geq 0\},\qquad q_{j}^{b}(x)=\mathbf{1}\{s_{q_{j}^{b}}(x)\geq 0\},$
with $q^{b}(x,j)=q_{j}^{b}(x)$.
B.4.1 Proof of Proposition 10
Proposition 10 (Natural separated construction).
For a realization $(x,a,y)$, the binary executed-pair loss admits the exact expansion
$\ell_{\mathrm{def\text{-}adv}}^{b}(r^{b},q^{b};\,x,a,y,\mathbf{e}):=\sum_{j=1}^{2}\sum_{k=0}^{1}c_{j,k}(x,a,y)\,\mathbf{1}\{r^{b}(x)=j\}\,\mathbf{1}\{q_{j}^{b}(x)=k\}.$ (18)
Replacing each indicator in (18) by its logistic upper bound yields the basic separated surrogate $\Phi_{\mathrm{def\text{-}adv}}^{b,r/q}(s_{r}^{b},\mathbf{s}_{q^{b}})$.
Proof This proposition records the most direct surrogate construction suggested by the protocol itself. In the binary setting there are only four executed outcomes: route to expert 1 or expert 2, and for the selected expert either do not query or query the single advice source. Writing these four cases out explicitly makes it clear how the usual logistic envelope enters.
Fix $(x,a,y)$. Since exactly one expert and one binary advice decision are executed, expanding the two sums in (18) gives
$\ell_{\mathrm{def\text{-}adv}}^{b}(r^{b},q^{b};\,x,a,y,\mathbf{e})$ $=c_{1,0}(x,a,y)\,\mathbf{1}\{s_{r}^{b}(x)<0\}\,\mathbf{1}\{s_{q_{1}^{b}}(x)<0\}$
$\quad+c_{1,1}(x,a,y)\,\mathbf{1}\{s_{r}^{b}(x)<0\}\,\mathbf{1}\{s_{q_{1}^{b}}(x)\geq 0\}$
$\quad+c_{2,0}(x,a,y)\,\mathbf{1}\{s_{r}^{b}(x)\geq 0\}\,\mathbf{1}\{s_{q_{2}^{b}}(x)<0\}$
$\quad+c_{2,1}(x,a,y)\,\mathbf{1}\{s_{r}^{b}(x)\geq 0\}\,\mathbf{1}\{s_{q_{2}^{b}}(x)\geq 0\}.$ (19)
This is the exact executed-pair loss written in terms of the decoded binary scores.
We now upper-bound each indicator by its logistic envelope [Steinwart, 2007, Bartlett et al., 2006]:
$\mathbf{1}\{s_{r}^{b}(x)<0\}\leq\Phi_{0}(s_{r}^{b}(x)),\qquad\mathbf{1}\{s_{r}^{b}(x)\geq 0\}\leq\Phi_{1}(s_{r}^{b}(x)),$
$\mathbf{1}\{s_{q_{j}^{b}}(x)<0\}\leq\Phi_{0}(s_{q_{j}^{b}}(x)),\qquad\mathbf{1}\{s_{q_{j}^{b}}(x)\geq 0\}\leq\Phi_{1}(s_{q_{j}^{b}}(x)),\qquad j\in\{1,2\}.$
Applying these four bounds term by term to (19) yields
$\ell_{\mathrm{def\text{-}adv}}^{b}(r^{b},q^{b};\,x,a,y,\mathbf{e})$ $\leq c_{1,0}(x,a,y)\,\Phi_{0}(s_{r}^{b}(x))\,\Phi_{0}(s_{q_{1}^{b}}(x))$
$\quad+c_{1,1}(x,a,y)\,\Phi_{0}(s_{r}^{b}(x))\,\Phi_{1}(s_{q_{1}^{b}}(x))$
$\quad+c_{2,0}(x,a,y)\,\Phi_{1}(s_{r}^{b}(x))\,\Phi_{0}(s_{q_{2}^{b}}(x))$
$\quad+c_{2,1}(x,a,y)\,\Phi_{1}(s_{r}^{b}(x))\,\Phi_{1}(s_{q_{2}^{b}}(x)).$
Grouping the first two terms and the last two terms by expert gives
$\ell_{\mathrm{def\text{-}adv}}^{b}(r^{b},q^{b};\,x,a,y,\mathbf{e})\leq\sum_{j=1}^{2}\Phi_{j-1}(s_{r}^{b}(x))\Bigl[c_{j,0}(x,a,y)\,\Phi_{0}(s_{q_{j}^{b}}(x))+c_{j,1}(x,a,y)\,\Phi_{1}(s_{q_{j}^{b}}(x))\Bigr].$
The right-hand side is exactly $\Phi_{\mathrm{def\text{-}adv}}^{b,r/q}(s_{r}^{b},\mathbf{s}_{q^{b}})$, namely the basic logistic-envelope construction described in Section 4.4. This proves the natural separated construction.
Regularity conditions for the larger separated family.
Equation (10) extends the basic construction by allowing monotone cost transforms, query reparameterizations, and flexible router weights. We impose the following conditions. They are not designed to make the counterexample artificially strong. They simply capture the regularity one would naturally ask of a well-behaved separated surrogate.
  1. (N1)
    Cost transform: $\nu:[0,C]\to(0,\infty)$ is continuous, strictly positive, differentiable on $(0,C)$ with $\nu^{\prime}(c)>0$.
  1. (Q1)
    Amplitude and reparameterization: $G:\mathbb{R}\to(0,\infty)$ is continuous and strictly positive; $U:\mathbb{R}\to\mathbb{R}$ is continuous, strictly increasing, and surjective.
  2. (Q2)
    Well-posedness: for every $(u,v)\in[0,C]^{2}$ not both zero, $t\mapsto\nu(u)\,G(t)\,\Phi_{0}(U(t))+\nu(v)\,G(t)\,\Phi_{1}(U(t))$ is coercive and strictly convex with a unique minimizer $t^{\star}(u,v)$.
  1. (P1)
    Router weights: $\Psi_{1},\Psi_{2}\in C^{2}(\mathbb{R})$ with $\Psi_{j}>0$, $\Psi_{1}^{\prime}<0$, $\Psi_{2}^{\prime}>0$, $\Psi_{j}^{\prime\prime}>0$.
  2. (P2)
    Routing ratio: $\rho(r)\coloneqq-\Psi_{1}^{\prime}(r)/\Psi_{2}^{\prime}(r)$ is a strictly decreasing bijection $\mathbb{R}\to(0,+\infty)$ with $\rho(0)=1$.
These conditions are satisfied, for instance, by the standard logistic (cross-entropy) surrogate with softmax router weights.
Containment of the traditional surrogate.
The familiar construction from the main text is recovered by taking $\nu(c)=c$, $G(q)\equiv 1$, and $U(q)=q$, together with the usual logistic router weights. The theorem proved below therefore applies to a broad family that still preserves the same separated router/query architecture. The next lemmas isolate the precise object this architecture optimizes.
The expert summary.
The next three lemmas formalize the quantity introduced in the main text. They show that after profiling over the query score of expert $j$, the entire row $(c_{j,0}(x,a,y),c_{j,1}(x,a,y))$ is compressed into the scalar summary $F(c_{j,0}(x,a,y),c_{j,1}(x,a,y))$. The key point is that this summary remains strictly sensitive to the non-minimal entry of the row. This is precisely where the surrogate departs from the Bayes comparison.
Definition 11 (Expert summary).
For $(u,v)\in[0,C]^{2}$, define
$F(u,v)\;\coloneqq\;\inf_{t\in\mathbb{R}}\;\bigl[\nu(u)\,G(t)\,\Phi_{0}(U(t))+\nu(v)\,G(t)\,\Phi_{1}(U(t))\bigr].$ (20)
B.4.2 Proof of Lemma 12
Lemma 12 (Well-defined expert summary).
Under (N1)(Q2), for every $(u,v)\in[0,C]^{2}$ not both zero, the infimum in (20) is attained at a unique point $t^{\star}(u,v)\in\mathbb{R}$.
Proof Fix $(u,v)\in[0,C]^{2}\setminus\{(0,0)\}$ and define
$f(t;u,v):=\nu(u)\,G(t)\,\Phi_{0}(U(t))+\nu(v)\,G(t)\,\Phi_{1}(U(t)).$
By (Q2), $f(\cdot;u,v)$ is coercive and strictly convex on $\mathbb{R}$. Coercivity implies that every sublevel set is compact, hence a global minimizer exists. Strict convexity then implies that this minimizer is unique.
B.4.3 Proof of Lemma 13
Lemma 13 (Expert summary after profiling).
Under (P1) and (Q2), for every fixed $s_{r}^{b}:\mathcal{X}\to\mathbb{R}$ and every realization $(x,a,y)$,
$\inf_{s_{q_{1}^{b}},\,s_{q_{2}^{b}}}\Phi_{\mathrm{def\text{-}adv}}^{\nu,r/q}(s_{r}^{b},\mathbf{s}_{q^{b}})$ $=\Psi_{1}(s_{r}^{b}(x))\,F(c_{1,0}(x,a,y),c_{1,1}(x,a,y))$ (21)
$+\Psi_{2}(s_{r}^{b}(x))\,F(c_{2,0}(x,a,y),c_{2,1}(x,a,y)).$
The profiled query minimizers are $t_{j}^{\star}=t^{\star}(c_{j,0}(x,a,y),c_{j,1}(x,a,y))$, hence they are independent of the router score $s_{r}^{b}$.
Proof Fix $(x,a,y)$ and write $t_{r}=s_{r}^{b}(x)$, $t_{1}=s_{q_{1}^{b}}(x)$, and $t_{2}=s_{q_{2}^{b}}(x)$. By (10),
$\Phi_{\mathrm{def\text{-}adv}}^{\nu,r/q}(s_{r}^{b},\mathbf{s}_{q^{b}})$ $=\Psi_{1}(t_{r})\Bigl[\nu(c_{1,0}(x,a,y))\,G(t_{1})\,\Phi_{0}(U(t_{1}))$
$\hskip 46.00012pt+\nu(c_{1,1}(x,a,y))\,G(t_{1})\,\Phi_{1}(U(t_{1}))\Bigr]$
$\quad+\Psi_{2}(t_{r})\Bigl[\nu(c_{2,0}(x,a,y))\,G(t_{2})\,\Phi_{0}(U(t_{2}))$
$\hskip 46.00012pt+\nu(c_{2,1}(x,a,y))\,G(t_{2})\,\Phi_{1}(U(t_{2}))\Bigr].$
The score $t_{1}$ appears only in the first bracket and $t_{2}$ appears only in the second. Because $\Psi_{1}(t_{r})$ and $\Psi_{2}(t_{r})$ are strictly positive by (P1), they factor through the infima over $t_{1}$ and $t_{2}$. Therefore,
$\inf_{s_{q_{1}^{b}},\,s_{q_{2}^{b}}}\Phi_{\mathrm{def\text{-}adv}}^{\nu,r/q}(s_{r}^{b},\mathbf{s}_{q^{b}})$
$=\Psi_{1}(t_{r})\,\inf_{t_{1}\in\mathbb{R}}\Bigl[\nu(c_{1,0}(x,a,y))\,G(t_{1})\,\Phi_{0}(U(t_{1}))+\nu(c_{1,1}(x,a,y))\,G(t_{1})\,\Phi_{1}(U(t_{1}))\Bigr]$
$\quad+\Psi_{2}(t_{r})\,\inf_{t_{2}\in\mathbb{R}}\Bigl[\nu(c_{2,0}(x,a,y))\,G(t_{2})\,\Phi_{0}(U(t_{2}))+\nu(c_{2,1}(x,a,y))\,G(t_{2})\,\Phi_{1}(U(t_{2}))\Bigr].$
Each infimum is exactly an instance of the expert summary (20). This gives (21). The characterization of the minimizers follows directly from Lemma 12.
B.4.4 Proof of Lemma 14
Lemma 14 (Continuity and strict monotonicity of the expert summary).
Under (N1)(Q2), the map $F:[0,C]^{2}\to(0,\infty)$ is continuous. Moreover, for every fixed $u\in(0,C)$, the map $v\mapsto F(u,v)$ is differentiable on $(0,C)$ and satisfies
$\frac{\partial F}{\partial v}(u,v)=\nu^{\prime}(v)\,G(t^{\star}(u,v))\,\Phi_{1}(U(t^{\star}(u,v)))>0.$ (22)
In particular, for every fixed $u\in(0,C)$, the function $v\mapsto F(u,v)$ is strictly increasing on $(0,C)$.
Proof Define
$f(t;u,v):=\nu(u)\,G(t)\,\Phi_{0}(U(t))+\nu(v)\,G(t)\,\Phi_{1}(U(t)).$
By (N1) and (Q1), the map $(t,u,v)\mapsto f(t;u,v)$ is continuous. By Lemma 12, each nonzero pair $(u,v)$ has a unique minimizer $t^{\star}(u,v)$. Standard continuity results for strictly convex coercive parametric minimization therefore imply that the map $(u,v)\mapsto t^{\star}(u,v)$ is continuous, and hence
$F(u,v)=f(t^{\star}(u,v);u,v)$
is continuous.
Now fix $u\in(0,C)$ and $v\in(0,C)$. Since the minimizer is unique, Danskin’s theorem applies to the function $v\mapsto\inf_{t\in\mathbb{R}}f(t;u,v)=F(u,v)$ and yields
$\frac{\partial F}{\partial v}(u,v)=\frac{\partial}{\partial v}f(t;u,v)\Big|_{t=t^{\star}(u,v)}.$
Differentiating $f$ with respect to $v$ gives
$\frac{\partial}{\partial v}f(t;u,v)=\nu^{\prime}(v)\,G(t)\,\Phi_{1}(U(t)).$
Substituting $t=t^{\star}(u,v)$ proves (22). Every factor on the right-hand side is strictly positive: $\nu^{\prime}(v)>0$ by (N1), $G(t^{\star}(u,v))>0$ by (Q1), and $\Phi_{1}(z)=\log(1+e^{z})>0$ for every $z\in\mathbb{R}$. Therefore $\partial F/\partial v(u,v)>0$, so $F(u,\cdot)$ is strictly increasing.
Why the expert summary creates the mismatch.
Lemmas 13 and 14 are the structural core of the impossibility result. Once the query scores are profiled out, expert $j$ no longer appears through its row minimum. It appears through the scalar summary $F(c_{j,0}(x,a,y),c_{j,1}(x,a,y))$, and this summary depends strictly on the non-minimal entry. Bayes and the surrogate are therefore comparing different objects.
B.4.5 Proof of Lemma 15
Lemma 15 (Router minimizer).
Under (P1)(P2), for any $A,B>0$ the function $t\mapsto A\,\Psi_{1}(t)+B\,\Psi_{2}(t)$ has a unique minimizer $t_{r}^{\star}(A,B)$ satisfying
$A>B\;\Longrightarrow\;t_{r}^{\star}>0,\qquad A=B\;\Longrightarrow\;t_{r}^{\star}=0,\qquad A<B\;\Longrightarrow\;t_{r}^{\star}<0.$
That is, the router selects the expert with the smaller profiled summary.
Proof Strict convexity follows from (P1), since
$\frac{d^{2}}{dt^{2}}\bigl[A\,\Psi_{1}(t)+B\,\Psi_{2}(t)\bigr]=A\,\Psi_{1}^{\prime\prime}(t)+B\,\Psi_{2}^{\prime\prime}(t)>0.$
Hence there is at most one minimizer. Differentiating the objective and setting the derivative to zero yields
$A\,\Psi_{1}^{\prime}(t_{r}^{\star})+B\,\Psi_{2}^{\prime}(t_{r}^{\star})=0,$
or equivalently
$\frac{\Psi_{1}^{\prime}(t_{r}^{\star})}{\Psi_{2}^{\prime}(t_{r}^{\star})}=\frac{B}{A}.$
The left-hand side is $\rho(t_{r}^{\star})$, which is a strictly decreasing bijection from $\mathbb{R}$ to $(0,\infty)$ by (P2). Since $\rho(0)=1$, we obtain
$\frac{B}{A}<1\iff t_{r}^{\star}(A,B)>0,\qquad\frac{B}{A}=1\iff t_{r}^{\star}(A,B)=0,\qquad\frac{B}{A}>1\iff t_{r}^{\star}(A,B)<0.$
This is exactly the stated sign characterization.
B.4.6 Completion of the proof of Theorem 6
See 6
Proof Fix $b\in(0,C)$ and $\varepsilon\in(0,C-b)$. We will construct a bounded cost table on which Bayes and the separated surrogate rank the two experts in opposite orders. The reason this can happen is the one identified in the main text: Bayes compares row minima, whereas the separated surrogate compares profiled row summaries.
Because $C>b+\varepsilon$ and $v\mapsto F(b,v)$ is strictly increasing by Lemma 14,
$F(b,C)>F(b,b+\varepsilon).$ (23)
The profile of the first row is therefore already larger than that of the second row when both rows have the same minimum value $b$. Since $(u,v)\mapsto F(u,v)$ is continuous, we can perturb the first row minimum slightly downward without destroying this strict inequality. More precisely, there exists $\delta\in(0,b)$ such that
$F(b-\delta,C)>F(b,b+\varepsilon).$ (24)
Now consider the cost table (12). Its row minima are
$\min\{b-\delta,C\}=b-\delta,\qquad\min\{b,b+\varepsilon\}=b.$
Since $\delta>0$, Bayes compares these two values and selects expert $1$.
The separated surrogate behaves differently. By Lemma 13, profiling out the two query scores reduces the pointwise objective to
$\widetilde{\Phi}_{\mathrm{def\text{-}adv}}^{\nu,r/q}(t)=F(b-\delta,C)\,\Psi_{1}(t)+F(b,b+\varepsilon)\,\Psi_{2}(t).$
Define
$A:=F(b-\delta,C),\qquad B:=F(b,b+\varepsilon).$
By (24), we have $A>B$. Lemma 15 then implies that the profiled router objective has a unique minimizer $t_{r}^{\star}(A,B)>0$. Under the decoding rule
$r^{b}(x)=1+\mathbf{1}\{s_{r}^{b}(x)\geq 0\},$
this router minimizer decodes to expert $2$.
To conclude, note that the query minimizers are also unique: for each row, Lemma 12 gives a unique minimizer $t^{\star}(c_{j,0},c_{j,1})$, and Lemma 13 shows that these minimizers are independent of the router score. Hence the full pointwise surrogate has a unique minimizer
$\bigl(t_{r}^{\star}(A,B),\,t^{\star}(b-\delta,C),\,t^{\star}(b,b+\varepsilon)\bigr),$
and this minimizer decodes to expert $2$ even though Bayes selects expert $1$. Thus the separated surrogate fails the defining pointwise criterion for Fisher consistency.
Interpretation.
The counterexample already appears within a bounded cost table, so the failure does not rely on any singular limiting construction. Nor does it disappear once the surrogate family is made more flexible: the contradiction persists under monotone cost transforms, positive query amplitudes, smooth reparameterizations, and well-behaved router weights. The obstruction is more basic. The separated architecture compresses each row before routing, whereas the Bayes rule compares rows only through their minima.
B.5 Proof of the Sequential–Composite Equivalence
Proposition 16 (Sequential–composite equivalence).
For every sequential policy $(r,q)$, the composite predictor $\pi(x)=(r(x),q(x,r(x)))$ incurs the same realized loss as $(r,q)$ under $\ell_{\mathrm{def\text{-}adv}}$. Conversely, every composite predictor $\pi:\mathcal{X}\to\Pi$ can be implemented by a sequential policy with the same realized loss. Consequently, the sequential and composite formulations induce the same population risks and therefore share the same Bayes-optimal decisions.
Proof The proof is simply an identification of the object that the loss depends on. Although the protocol is written sequentially, the realized loss only sees the executed pair consisting of the routed expert and the advice ultimately shown to that expert.
Let $(r,q)$ be any sequential policy and define $\pi(x)=(r(x),q(x,r(x)))$. By Definition 2,
$\ell_{\mathrm{def\text{-}adv}}(r,q;\,x,a,y,\mathbf{e})=c_{r(x),\,q(x,r(x))}(x,a,y)=c_{\pi_{1}(x),\,\pi_{2}(x)}(x,a,y)=\ell_{\mathrm{def\text{-}adv}}(\pi;\,x,a,y,\mathbf{e}),$
where $\pi(x)=(\pi_{1}(x),\pi_{2}(x))$. Thus every sequential policy induces a composite predictor with exactly the same realized loss at every input.
Conversely, let $\pi:\mathcal{X}\to\Pi$ and write $\pi(x)=(\pi_{1}(x),\pi_{2}(x))$. To implement the same executed pair sequentially, define
$r(x)=\pi_{1}(x).$
The query function only matters at the routed expert, so it is enough to require
$q(x,r(x))=\pi_{2}(x).$
One convenient choice is to set $q(x,j)=\pi_{2}(x)$ for every $j\in[J]$. Then $q(x,r(x))=\pi_{2}(x)$ and therefore
$\ell_{\mathrm{def\text{-}adv}}(r,q;\,x,a,y,\mathbf{e})=c_{\pi_{1}(x),\,\pi_{2}(x)}(x,a,y)=\ell_{\mathrm{def\text{-}adv}}(\pi;\,x,a,y,\mathbf{e}).$
So every composite predictor can be realized by a sequential policy with the same pointwise loss.
Since the two constructions agree pointwise in both directions, the sequential and composite formulations induce exactly the same set of achievable realized losses. Their population risk functionals therefore have the same infima and the same Bayes-optimal decisions.
B.6 Proof of Proposition 17
Proposition 17 (Mismatch decomposition).
For every composite predictor $\pi:\mathcal{X}\to\Pi$ and every realization $(x,a,y)$,
$\ell_{\mathrm{def\text{-}adv}}(\pi;\,x,a,y,\mathbf{e})\;=\;D(x,a,y)\;+\;\sum_{i\in\Pi}w_{i}(x,a,y)\,\mathbf{1}\{\pi(x)\neq i\},$
where $M(x,a,y)=\max_{t\in\Pi}c_{t}(x,a,y)$,  $w_{i}(x,a,y)=M(x,a,y)-c_{i}(x,a,y)\geq 0$,  and $D(x,a,y)=\sum_{i\in\Pi}c_{i}(x,a,y)-(|\Pi|-1)\,M(x,a,y)$.
Proof The decomposition is easiest to see by fixing the composite action that is actually executed. Let
$i_{\pi}\coloneqq\pi(x)\in\Pi.$
Since the predictor selects exactly one composite action, the true loss is just the realized cost of that action:
$\ell_{\mathrm{def\text{-}adv}}(\pi;\,x,a,y,\mathbf{e})\;=\;c_{i_{\pi}}(x,a,y).$
Now expand the mismatch term. The indicator $\mathbf{1}\{\pi(x)\neq i\}$ vanishes only at the executed action $i=i_{\pi}$, so the sum keeps every weight except the one attached to the chosen pair:
$\sum_{i\in\Pi}w_{i}\,\mathbf{1}\{\pi(x)\neq i\}$ $=\sum_{i\neq i_{\pi}}w_{i}=\sum_{i\neq i_{\pi}}\bigl(M-c_{i}\bigr)$
$=(|\Pi|-1)\,M-\sum_{i\neq i_{\pi}}c_{i}.$
Adding the offset
$D=\sum_{i\in\Pi}c_{i}-(|\Pi|-1)\,M$
therefore gives
$D+\sum_{i\in\Pi}w_{i}\,\mathbf{1}\{\pi(x)\neq i\}=\sum_{i\in\Pi}c_{i}-(|\Pi|-1)\,M+(|\Pi|-1)\,M-\sum_{i\neq i_{\pi}}c_{i}=c_{i_{\pi}},$
which is exactly $\ell_{\mathrm{def\text{-}adv}}(\pi;\,x,a,y,\mathbf{e})$. This proves the decomposition.
B.7 Proof of Lemma 7
See 7
Proof This lemma is a derivation rather than a separate statistical claim. Once the true loss is written over the composite action space, the only action-dependent part is the weighted mismatch term. The augmented surrogate is obtained by replacing each mismatch indicator with a multiclass surrogate on $\Pi$.
Indeed, Proposition 17 shows that
$\ell_{\mathrm{def\text{-}adv}}(\pi;\,x,a,y,\mathbf{e})\;=\;D(x,a,y)+\sum_{i\in\Pi}w_{i}(x,a,y)\,\mathbf{1}\{\pi(x)\neq i\},$
where $D(x,a,y)$ does not depend on the chosen action. Thus, for surrogate design, the relevant object is
$\sum_{i\in\Pi}w_{i}(x,a,y)\,\mathbf{1}\{\pi(x)\neq i\}.$
The comp-sum family (14) provides a multiclass surrogate for prediction on the composite label space $\Pi$. Applying it label by label to the weighted mismatch term leads directly to
$\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}(s_{\pi};\,x,a,y,\mathbf{e})\;\coloneqq\;\sum_{i\in\Pi}w_{i}(x,a,y)\,\Phi^{\tau}_{01}(\mathbf{s}_{\pi}(x),i),$
which is exactly (15). In other words, the augmented surrogate is the comp-sum analogue of the weighted mismatch representation of the true loss on the composite action space.
B.8 Proof of Theorem 8
The proof proceeds by reducing the augmented objective to an ordinary multiclass problem on the composite action space $\Pi$, with example-dependent weights absorbed into a new data distribution.
B.8.1 Proof of Lemma 18
Lemma 18 (Weighted multiclass reduction).
Let $\mathbf{w}(x)\coloneqq(\mathbb{E}[w_{i}(X,A,Y)\mid X=x])_{i\in\Pi}$, and define
$d(x)\coloneqq\mathbb{E}[D(X,A,Y)\mid X=x].$
For every $i\in\Pi$, write
$\bar{w}_{i}(x)\coloneqq\mathbb{E}[w_{i}(X,A,Y)\mid X=x].$
Define a probability vector $p(x)=(p_{i}(x))_{i\in\Pi}$ on $\Pi$ by
$p_{i}(x)\coloneqq\begin{cases}\bar{w}_{i}(x)/\|\mathbf{w}(x)\|_{1},&\|\mathbf{w}(x)\|_{1}>0,\\ 1/|\Pi|,&\|\mathbf{w}(x)\|_{1}=0,\end{cases}\qquad i\in\Pi,$
and, for any composite predictor $\pi$ and any score function $s_{\pi}$, define the conditional multiclass risks
$C_{01}^{p}(\pi,x)\coloneqq\sum_{i\in\Pi}p_{i}(x)\,\mathbf{1}\{\pi(x)\neq i\},\qquad C_{\Phi^{\tau}_{01}}^{p}(s_{\pi},x)\coloneqq\sum_{i\in\Pi}p_{i}(x)\,\Phi^{\tau}_{01}(\mathbf{s}_{\pi}(x),i).$
Assume $\mathbb{E}[\|\mathbf{w}(X)\|_{1}]>0$, and define a probability distribution $\widetilde{\mathcal{D}}$ on $\mathcal{X}\times\Pi$ by
$\widetilde{\mathbb{P}}(X\in B,I=i)\;\coloneqq\;\frac{1}{\mathbb{E}[\|\mathbf{w}(X)\|_{1}]}\,\mathbb{E}\bigl[\bar{w}_{i}(X)\,\mathbf{1}\{X\in B\}\bigr]$ (25)
for every measurable set $B\subseteq\mathcal{X}$ and every $i\in\Pi$. Let $\ell_{01}(\pi;\,x,i)\coloneqq\mathbf{1}\{\pi(x)\neq i\}$ denote the multiclass 011 loss on $\Pi$. Then, for every scoring function $s_{\pi}\in\mathcal{H}_{\pi}$ and its decoded predictor $\pi(x)=\operatorname*{arg\,max}_{i\in\Pi}s_{\pi}(x,i)$,
$C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)$ $=d(x)+\|\mathbf{w}(x)\|_{1}\,C_{01}^{p}(\pi,x),$ (26)
$C_{\Phi^{\mathrm{aug},\tau}}(s_{\pi},x)$ $=\|\mathbf{w}(x)\|_{1}\,C_{\Phi^{\tau}_{01}}^{p}(s_{\pi},x),$ (27)
where
$C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)\coloneqq\mathbb{E}\bigl[\ell_{\mathrm{def\text{-}adv}}(\pi;\,X,A,Y,\mathbf{e})\mid X=x\bigr]$
and
$C_{\Phi^{\mathrm{aug},\tau}}(s_{\pi},x)\coloneqq\mathbb{E}\bigl[\Phi_{\mathrm{def\text{-}adv}}^{\mathrm{aug},\tau}(s_{\pi};\,X,A,Y,\mathbf{e})\mid X=x\bigr].$
Consequently,
[…]
Moreover, the best-in-class risks and minimizability gaps satisfy
[…]
Proof The point of the lemma is that, once we condition on $X=x$, the deferral-advice problem becomes an ordinary multiclass problem on the composite action space $\Pi$, up to an additive offset and a positive scale.
Fix $x$. By Proposition 17,
$\ell_{\mathrm{def\text{-}adv}}(\pi;\,X,A,Y,\mathbf{e})=D(X,A,Y)+\sum_{i\in\Pi}w_{i}(X,A,Y)\,\mathbf{1}\{\pi(X)\neq i\}.$
Taking the conditional expectation given $X=x$ gives
$C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)=d(x)+\sum_{i\in\Pi}\bar{w}_{i}(x)\,\mathbf{1}\{\pi(x)\neq i\}.$
If $\|\mathbf{w}(x)\|_{1}=0$, then every $\bar{w}_{i}(x)$ vanishes, so the action-dependent term disappears and
$C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)=d(x).$
Because the multiplier $\|\mathbf{w}(x)\|_{1}$ is also zero, both (26) and (27) hold automatically in this degenerate case, regardless of the arbitrary choice of $p(x)$.
Assume now that $\|\mathbf{w}(x)\|_{1}>0$. Then $\bar{w}_{i}(x)=\|\mathbf{w}(x)\|_{1}\,p_{i}(x)$ for every $i\in\Pi$, and therefore
$C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)=d(x)+\|\mathbf{w}(x)\|_{1}\sum_{i\in\Pi}p_{i}(x)\,\mathbf{1}\{\pi(x)\neq i\},$
which is exactly (26).
The same normalization applies to the surrogate:
$C_{\Phi^{\mathrm{aug},\tau}}(s_{\pi},x)=\sum_{i\in\Pi}\bar{w}_{i}(x)\,\Phi^{\tau}_{01}(\mathbf{s}_{\pi}(x),i)=\|\mathbf{w}(x)\|_{1}\sum_{i\in\Pi}p_{i}(x)\,\Phi^{\tau}_{01}(\mathbf{s}_{\pi}(x),i),$
which proves (27).
Thus, for each fixed input $x$, the conditional deferral-advice problem is exactly a weighted multiclass problem on $\Pi$, up to the action-independent offset $d(x)$.
To pass from this conditional reduction to a population statement, we average these weighted multiclass problems over $X$. Since $\bar{w}_{i}(X)\geq 0$ for all $i\in\Pi$, the right-hand side of (25) defines a finite non-negative measure on $\mathcal{X}\times\Pi$. Its total mass is
$\sum_{i\in\Pi}\widetilde{\mathbb{P}}(X\in\mathcal{X},I=i)=\frac{1}{\mathbb{E}[\|\mathbf{w}(X)\|_{1}]}\sum_{i\in\Pi}\mathbb{E}[\bar{w}_{i}(X)]=\frac{1}{\mathbb{E}[\|\mathbf{w}(X)\|_{1}]}\,\mathbb{E}\bigl[\|\mathbf{w}(X)\|_{1}\bigr]=1,$
so $\widetilde{\mathcal{D}}$ is a probability distribution.
Integrating (27) over $X$ yields
[…]
which is (28). Likewise, (26) gives
[…]
which is (29).
The remaining identities follow from the same affine scaling. Taking infima over $s_{\pi}\in\mathcal{H}_{\pi}$ or over the induced predictors gives (30) and (32). Since the minimizability gaps are defined through these same best-in-class risks, the same argument also yields (31) and (33).
See 8
Proof If $\mathbb{E}[\|\mathbf{w}(X)\|_{1}]=0$, then $\bar{w}_{i}(X)=0$ almost surely for every $i\in\Pi$. By Proposition 17, the true risk then reduces to the expectation of the action-independent term $D(X,A,Y)$, so every predictor is optimal and the theorem is trivial. We therefore assume $\mathbb{E}[\|\mathbf{w}(X)\|_{1}]>0$.
The key idea is to compare the surrogate and the true loss at the level where the Bayes policy itself is defined, namely conditionally on the input $x$. Lemma 18 shows that, once $X=x$ is fixed, both the true deferral-advice risk and the augmented surrogate reduce to weighted multiclass risks on the composite action space $\Pi$:
$C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)=d(x)+\|\mathbf{w}(x)\|_{1}\,C_{01}^{p}(\pi,x),\qquad C_{\Phi^{\mathrm{aug},\tau}}(s_{\pi},x)=\|\mathbf{w}(x)\|_{1}\,C_{\Phi^{\tau}_{01}}^{p}(s_{\pi},x).$
The additive term $d(x)$ is independent of the chosen composite action, so it plays no role in the comparison between policies. What matters is that the action-dependent part of the true loss and the surrogate are governed by the same conditional weight vector $p(x)$.
It is therefore natural to look first at the conditional excess risk. For each input $x$, define
$\Delta C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)$ $\coloneqq C_{\ell_{\mathrm{def\text{-}adv}}}(\pi,x)-\inf_{\pi^{\prime}\in\mathcal{H}_{\pi}}C_{\ell_{\mathrm{def\text{-}adv}}}(\pi^{\prime},x),$
$\Delta C_{\Phi^{\mathrm{aug},\tau}}(s_{\pi},x)$ $\coloneqq C_{\Phi^{\mathrm{aug},\tau}}(s_{\pi},x)-\inf_{s^{\prime}_{\pi}\in\mathcal{H}_{\pi}}C_{\Phi^{\mathrm{aug},\tau}}(s^{\prime}_{\pi},x).$
When $\|\mathbf{w}(x)\|_{1}>0$, the identities above show that these conditional gaps are exactly the weighted multiclass excess risks induced by $p(x)$, up to the common scale factor $\|\mathbf{w}(x)\|_{1}$. This is the local reason the augmented surrogate is aligned with the correct comparison. The remainder of the proof turns this pointwise observation into an $\mathcal{H}_{\pi}$-consistency statement at the population level.
This identifies the right conditional comparison, but the theorem is an $\mathcal{H}_{\pi}$-consistency statement rather than a purely pointwise calibration claim. To keep the hypothesis-class restriction explicit, we gather these conditional weighted problems into the multiclass distribution $\widetilde{\mathcal{D}}$ defined in Lemma 18. Under that distribution, the comp-sum loss $\Phi^{\tau}_{01}$ satisfies the usual multiclass $\mathcal{H}_{\pi}$-consistency inequality with transfer function $\Gamma_{\tau}$. Write
[…]
Then
[…]
Lemma 18 shows that this ordinary multiclass problem is exactly the population counterpart of the conditional reduction above. Define
$\widetilde{\Gamma}_{\tau}(u)\;\coloneqq\;\mathbb{E}\bigl[\|\mathbf{w}(X)\|_{1}\bigr]\,\Gamma_{\tau}\!\left(\frac{u}{\mathbb{E}[\|\mathbf{w}(X)\|_{1}]}\right).$
Multiplying by $\mathbb{E}[\|\mathbf{w}(X)\|_{1}]$ gives
[…]
We now translate this inequality back to deferral with advice using (28)–(33). The additive term $\mathbb{E}[D(X,A,Y)]$ appears in both the incurred true risk and the best-in-class true risk, so it cancels from the excess-risk difference. We obtain
$\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\pi)-\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}^{B}(\mathcal{H}_{\pi})+\mathcal{U}_{\ell_{\mathrm{def\text{-}adv}}}(\mathcal{H}_{\pi})\leq\widetilde{\Gamma}_{\tau}\!\left(\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}(s_{\pi})-\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}^{\ast}(\mathcal{H}_{\pi})+\mathcal{U}_{\Phi^{\mathrm{aug},\tau}}(\mathcal{H}_{\pi})\right),$
which is exactly (16).
This proves the claimed $\mathcal{H}_{\pi}$-consistency bound. When the minimizability gaps vanish, the right-hand side tends to zero whenever the surrogate excess risk tends to zero, so the augmented surrogate recovers the Bayes-optimal deferral-advice risk.
From Mao et al. [2023b], we have:
$\Gamma_{\tau}(u)=\begin{cases}\sqrt{2^{\tau}(2-\tau)u}&\tau\in[0,1)\\ \sqrt{2|\Pi|^{\tau-1}v})&\tau\in[1,2)\\ (\tau-1)|\Pi|^{\tau-1}v&\tau\in[2,+\infty)\end{cases}$ (34)
B.9 Proof of Corollary 9
Proof The corollary is the statistical form of Theorem 8. Once the two minimizability gaps vanish, the theorem compares the true excess risk of the induced policy directly to the surrogate excess risk of the learned score.
Under the stated assumption,
$\mathcal{U}_{\ell_{\mathrm{def\text{-}adv}}}(\mathcal{H}_{\pi})=\mathcal{U}_{\Phi^{\mathrm{aug},\tau}}(\mathcal{H}_{\pi})=0.$
Moreover, vanishing true minimizability gap implies
$\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}^{B}(\mathcal{H}_{\pi})=\inf_{\pi}\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\pi).$
Applying Theorem 8 to the learned score $\widehat{s}_{\pi,n}$ and its induced predictor $\widehat{\pi}_{n}$ therefore gives, for every $n$,
$\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\widehat{\pi}_{n})-\inf_{\pi}\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\pi)\;\leq\;\widetilde{\Gamma}_{\tau}\!\left(\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}(\widehat{s}_{\pi,n})-\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}^{\ast}(\mathcal{H}_{\pi})\right).$
Thus the only remaining quantity to control is the surrogate excess risk.
By assumption,
$\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}(\widehat{s}_{\pi,n})-\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}^{\ast}(\mathcal{H}_{\pi})\to 0$
in probability. Since $\widetilde{\Gamma}_{\tau}$ is non-negative, non-decreasing, and satisfies $\widetilde{\Gamma}_{\tau}(0)=0$, it follows that
$\widetilde{\Gamma}_{\tau}\!\left(\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}(\widehat{s}_{\pi,n})-\mathcal{E}_{\Phi^{\mathrm{aug},\tau}}^{\ast}(\mathcal{H}_{\pi})\right)\to 0$
in probability as well. Combining this with the preceding excess-risk bound proves that
$\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\widehat{\pi}_{n})-\inf_{\pi}\mathcal{E}_{\ell_{\mathrm{def\text{-}adv}}}(\pi)\to 0$
in probability.
Appendix C Experiments
C.1 Global Experimental Protocol
Before turning to the individual tasks, we summarize the experimental choices that are shared across the appendix. The goal is to evaluate the learned policy in the regime studied by the theory. The learner does not predict an abstract label directly. It chooses an executed expert–advice pair, and the value of that choice is revealed only through the resulting realized cost. The experimental protocol is therefore designed to mirror the formal setup as closely as possible.
Evaluation metric.
Our primary metric is always the empirical average of the true deferral-advice loss,
$\widehat{\mathcal{E}}_{\ell_{\mathrm{def\text{-}adv}}}(\pi)\;=\;\frac{1}{n_{\mathrm{val}}}\sum_{i=1}^{n_{\mathrm{val}}}\ell_{\mathrm{def\text{-}adv}}(\pi;x_{i},a_{i},y_{i},\mathbf{e}_{i}),$
evaluated on the validation split of the corresponding experiment. This is the only quantity used to compare policies across values of $\lambda$. The reason is structural. Once advice costs vary, two policies may have very different consultation patterns even when they achieve similar predictive accuracy. Accuracy alone therefore no longer represents the deployment objective of the paper. The true loss remains the correct comparison criterion because it accounts jointly for task error, expert fees, and advice costs. The appendix therefore keeps the focus on true-loss tables and uses advice-distribution figures only as diagnostics to explain how a method achieves a certain true loss rather than to rank methods.
Learned policy.
In every experiment, the learned policy acts on the composite action space $\Pi=[J]\times[K]_{0}$. The deployed decision is therefore a single executed pair $(j,k)$. Still, it is useful to parameterize the scores in a way that mirrors the protocol. We write the score of action $(j,k)$ as
$s_{\theta}(x,(j,k))\;=\;u_{\theta}(x,j)+v_{\theta}(x,j,k),$
where $u_{\theta}(x,j)$ plays the role of an expert-routing score and $v_{\theta}(x,j,k)$ is an expert-conditional advice adjustment. Both depend on the input through a shared learned representation, so the same construction can be used for text and tabular inputs.
A convenient instantiation first maps the input to a latent representation $z_{\theta}(x)\in\mathbb{R}^{d}$ and then introduces routing embeddings $a_{j}\in\mathbb{R}^{d}$, advice-side expert embeddings $m_{j}\in\mathbb{R}^{d}$, advice embeddings $g_{k}\in\mathbb{R}^{d}$, and score biases $\rho_{j},\delta_{k}\in\mathbb{R}$:
$u_{\theta}(x,j)$ $:=\rho_{j}+\langle a_{j},\;z_{\theta}(x)\rangle,$
$v_{\theta}(x,j,k)$ $:=\delta_{k}+\big\langle m_{j},\;g_{k}\odot z_{\theta}(x)\big\rangle,$
where $\odot$ denotes elementwise multiplication. This factorization preserves three properties that are important in practice. First, routing remains input-dependent through $u_{\theta}$. Second, the value of advice acquisition can depend jointly on the input and an expert through $v_{\theta}$. Third, the expert–advice interaction is non-separable, while the number of parameters still scales through a shared low-rank representation rather than independently over all $J(K+1)$ composite actions.
Decoding remains a single argmax over composite actions:
$\pi_{\theta}(x)\;=\;\operatorname*{arg\,max}_{(j,k)\in\Pi}\;s_{\theta}(x,(j,k)).$
[…]
Figure 2: Structured parameterization of the composite policy score. The final decision remains a single argmax over executed expert–advice pairs, while the score decomposes into a routing term and an expert-conditional advice adjustment.
Resources.
All experiments were conducted on an internal cluster using an NVIDIA A100 GPU with 40 GB of VRAM.
Reading the appendix tables.
The main paper reports only the true deferral-advice loss, because that is the deployment metric of interest and the one that remains meaningful across cost regimes. The appendix keeps the same true-loss tables, but adds expert-by-advice summaries and advice-distribution figures for selected values of $\lambda$. These additional diagnostics help explain how the learned policy adapts as advice becomes more or less expensive, and they make explicit how the deployment-optimal expert-advice pair changes with the cost regime.
C.1.1 Baselines
We use the same baseline family throughout, and each baseline is included for a specific reason. All methods are evaluated on exactly the same precomputed executed-pair table, so differences in performance come from the decision rule rather than from stochastic variation in the experts or advice sources.
Our method. This is our main method. It predicts directly over the composite action space $\Pi=[J]\times[K]_{0}$ and is the only learned policy that is allowed to coordinate expert selection and advice acquisition jointly. Its role is therefore to measure the benefit of solving the full deferral-advice problem.
L2D. This is the standard no-query baseline obtained by restricting the action space to the $J$ actions $(j,0)$ recovering [Mao et al., 2023a, 2024c, Montreuil et al., 2025b]. It answers the first natural question: is explicit advice acquisition useful at all beyond ordinary learning to defer? If our method does not improve over L2D, then the added advice mechanism is not buying anything under the true deployment objective.
Best fixed pair. This baseline selects a single constant composite action $(j,k)$ and applies it to every input. It is an oracle over non-adaptive policies, and therefore measures how much is gained by input-dependent decision-making rather than by choosing one good expert–advice pair globally.
Partial-randomization ablations. These are the two most informative diagnostic baselines. In “learned expert, random advice,” we keep the expert chosen by our learned policy and replace only the advice action by a uniform random choice. In “random expert, learned advice,” we do the reverse. Taken together, these two baselines separate the contribution of expert selection from that of advice selection. If the full method improves over both, then the gain cannot be attributed to only one half of the decision.
Random $(j,k)$. The final baseline samples both the expert and the advice action uniformly. Its role is not diagnostic but calibrating: it provides an uninformed floor against which the learned and structured methods can be compared.
C.2 Synthetic Theorem-Aligned Experiment
This appendix includes one synthetic benchmark whose role is purely theoretical. The goal is not to mimic a realistic deployment pipeline, but to instantiate the exact mechanism of Theorem 6 and show that the observed failure of the separated surrogate is the one predicted by the theory. The same benchmark also contains a second region in which advice is genuinely useful, so that standard L2D is structurally suboptimal. The synthetic experiment therefore isolates the two distinct reasons why the augmented formulation should dominate the alternatives.
Goal, data, and metric.
We work in the minimal binary setting of the theorem: two experts ($J=2$) and one advice action ($K=1$), so the composite action space is $\Pi=\{(1,0),(1,1),(2,0),(2,1)\}$. Inputs are sampled uniformly from $1,1]^{2}$, and the first coordinate partitions the space into two regions:
$R_{-}\coloneqq\{x\in[-1,1]^{2}:x_{1}<0\},\qquad R_{+}\coloneqq\{x\in[-1,1]^{2}:x_{1}\geq 0\}.$
The second coordinate is irrelevant to the Bayes rule; it is included only so that the learned decision maps can be visualized on a two-dimensional grid. We evaluate every policy by its expected true deferral-advice loss on an independent test split drawn from the same synthetic environment. Because the environment is defined directly through executed costs, the evaluation metric matches the object studied in the theory exactly.
Certified cost construction.
We fix $\lambda_{0}=0.08$ and assign one conditional executed-cost table to each region. On $R_{-}$, we use the certified theorem table
$T^{-}=\begin{pmatrix}0.38&1.08\\ 0.50&0.51\end{pmatrix}$,
while on $R_{+}$ we use the advice-helpful table
$T^{+}=\begin{pmatrix}0.55&0.18\\ 0.30&0.90\end{pmatrix}$.
Table 4 records these two regions in the same executed-action format as the main paper. Every queried entry is realizable as a Bernoulli task loss plus the fixed advice fee $\lambda_{0}$, so the synthetic environment is not an abstract cost table detached from any prediction model.
Table 4: Certified conditional executed-cost tables used in the synthetic benchmark. Each row is one region of the input space, and each column is one executed expert–advice pair.
Region $(1,0)$ $(1,1)$ $(2,0)$ $(2,1)$ Bayes action Bayes risk
$R_{-}$ 0.380 1.080 0.500 0.510 $(1,0)$ 0.380
$R_{+}$ 0.550 0.180 0.300 0.900 $(1,1)$ 0.180
The left region is the theorem region. Bayes compares row minima and therefore prefers expert $1$, since $\min_{k}T^{-}_{1,k}=0.38<0.50=\min_{k}T^{-}_{2,k}$. The exact binary separated surrogate from the paper behaves differently: its profiled row summaries satisfy
F(0.38,1.08)=0.8371> 0.7000=F(0.50,0.51),F(0.38,1.08)=0.8371\;>\;0.7000=F(0.50,0.51),
so the profiled surrogate reverses the comparison and favors expert $2$. The right region serves a different purpose. There, Bayes chooses the queried action $(1,1)$ with cost $0.18$, whereas L2D is restricted to the no-advice column and must therefore pick expert $2$ with cost $0.30$. Since both regions have positive probability, the separated surrogate and L2D each incur a positive irreducible excess risk, but for different reasons.
Policies and exact separated baseline.
We compare our method, L2D, the same random baselines used in the real experiments, and one exact implementation of the binary separated surrogate analyzed in Section 4.4. This is important. The separated baseline is not an approximate router-query variant. It is precisely the special case obtained by setting $\nu=\mathrm{id}$, $G\equiv 1$, $U=\mathrm{id}$, and $\Psi_{j}=\Phi_{j-1}$, so that each indicator in the true binary loss is replaced by its logistic envelope. Concretely, the model uses one binary router score $s_{r}^{b}$ and one expert-conditional query score $s_{q_{1}^{b}}$ or $s_{q_{2}^{b}}$ per row, exactly as in the theorem statement.
Training protocol.
All learned baselines use the same two-layer MLP over the synthetic input $x\in\mathbb{R}^{2}$, with hidden widths $(32,32)$ and ReLU activations. The Our method is trained with the augmented comp-sum surrogate at $\tau=1$, while the separated baseline is trained with the exact binary logistic surrogate described above. We sweep the training size over $n\in\{250,500,1000,2500,5000\}$ and report averages over five seeds. The figures and tables below use the largest train size, $n=5000$, for the pointwise comparisons.
Table 5: Synthetic test performance at the largest train size ($n=5000$). Our method is the only one whose excess risk is essentially zero.
Method Test true risk Test excess risk Advice rate (%) Bayes-action match (%)
Our $281\pm 0.001 0.0$$001\pm 0.000$ 49.6 99.3
Separated $334\pm 0.004 0.0$$054\pm 0.004$ 62.5 58.4
L2D $342\pm 0.001 0.0$$62\pm 0.000 0.0$ 48.8
Learned route, random advice $0.661\pm 0.002$ $0.381\pm 0.002$ 49.6 24.5
Random route, no advice $0.432\pm 0.001$ $152\pm 0.001 0.0$ 24.9
Random $(j,k)$ $0.550\pm 0.003$ $0.269\pm 0.004$ 50.1 24.9
Table 5 is the central quantitative result. Our method reaches $0.281\pm 0.001$ true risk, essentially matching the Bayes optimum of $0.280$. The separated surrogate stalls far above this level, with excess risk $0.054\pm 0.004$, even though it is trained with the exact binary surrogate analyzed in the theorem. L2D also remains suboptimal, but for a different reason: it never queries, so it cannot realize the Bayes action on $R_{+}$. The random baselines are much worse, which confirms that the gain is not coming from a trivial bias in the environment.
Table 6: Regionwise diagnosis at the largest train size ($n=5000$). The theorem region $R_{-}$ isolates the failure of the separated surrogate, while the advice-helpful region $R_{+}$ isolates the limitation of L2D.
Method Risk on $R_{-}$ Risk on $R_{+}$ Bayes match on $R_{-}$ (%) Bayes match on $R_{+}$ (%)
Our 0.381 0.181 99.3 99.3
Separated 0.484 0.184 18.0 99.0
L2D 0.383 0.300 97.5 0.0
Learned route, random advice 0.727 0.594 48.9 0.1
Random route, no advice 0.440 0.424 49.8 0.0
Random $(j,k)$ 0.617 0.482 24.7 25.1
Table 6 shows that the two failure modes are exactly the ones predicted by the theory. On $R_{-}$, the separated surrogate matches the Bayes action only $18.0\%$ of the time, even though it is nearly perfect on $R_{+}$; this is the empirical footprint of the profiled-row distortion proved in Theorem 6. L2D exhibits the opposite pattern. It is nearly optimal on $R_{-}$, where the Bayes action is no-advice, but it matches the Bayes action $0\%$ of the time on $R_{+}$ because the Bayes decision there is the queried action $(1,1)$. Our method is nearly perfect on both regions, which is exactly the consistency story established by Theorem 8.
Figure 3: Synthetic test excess risk as a function of train size. Our method approaches zero excess risk, whereas the exact separated surrogate and L2D remain bounded away from the Bayes optimum.
Figure 4: Learned decision maps at the largest train size ($n=5000$). The vertical dashed line marks the transition between the theorem region $R_{-}$ and the advice-helpful region $R_{+}$. Our method recovers the Bayes split, the separated surrogate selects the wrong expert on the left region, and L2D cannot realize the queried Bayes action on the right region.
Figures 3 and 4 make the same point visually. The excess risk curve shows that our method is the only one that continues to improve toward Bayes optimality as the sample size grows. The action maps then show why: our method learns the correct composite decision on both halves of the space, while the exact separated surrogate and L2D fail on the region singled out by the corresponding theory.
C.3 FEVER Experimental Details
This appendix records the implementation details that are omitted from the main paper for space reasons. We focus on the FEVER [Thorne et al., 2018] experiment reported in Section 5, since it is the main empirical study in the paper.
Data, action space, and metric.
FEVER [Thorne et al., 2018] is a fact-verification benchmark built from short natural-language claims paired with Wikipedia evidence. Each example must be classified as SUPPORTS, REFUTES, or NOT_ENOUGH_INFO depending on whether the claim is supported by the evidence available in Wikipedia. A typical example is a claim such as “The Eiffel Tower is located in Berlin”; the correct label is REFUTES once the relevant Wikipedia passages are retrieved. The experiment uses 15,00015{,}000 examples sampled from the labelled development split, with a stratified $75/25$ train-validation partition. The composite action space has size $3\times 5$, corresponding to three experts and five advice actions. All policy comparisons use the global protocol of Section C.1; in particular, the reported numbers are validation averages of the true deferral-advice loss.
Experts and advice construction.
The expert pool contains Qwen3-4B-Instruct [Yang et al., 2025], Qwen3-8B, and DeBERTa-large [He et al., 2021] trained on claim tasks, with respective consultation costs $\beta=(0.03,0.05,0.04)$. Advice is derived from Wikipedia evidence retrieved with BM25 [Robertson and Zaragoza, 2009]. The five advice actions are: no retrieval, BM25 top-$1$, BM25 top-$3$, BM25 top-$5$, and query reformulation followed by BM25 top-$5$. The reformulated query is produced by Qwen2.5-1.5B. Query costs take the form $\gamma_{k}=\lambda\gamma_{k}^{\mathrm{base}}$ with $\gamma^{\mathrm{base}}=(0,0.015,0.02,0.03,0.01)$. This construction is chosen to induce heterogeneous expert-advice trade-offs, rather than to mimic exact monetary costs.
Prompt templates.
The expert-facing prompts are deliberately simple so that the experiment isolates the interaction between routing and advice rather than prompt engineering. For the two Qwen experts, the system instruction is to return exactly one of SUPPORTS, REFUTES, or NOT_ENOUGH_INFO. The user message then instantiates the template
Claim: <claim>.
Evidence: <retrieved passages, or EMPTY when k=0k=0>.
Answer with exactly one label: SUPPORTS, REFUTES, or NOT_ENOUGH_INFO.
For example, if the claim is “The Eiffel Tower is located in Berlin,” the top-$1$ advice may supply a passage stating that the Eiffel Tower is in Paris, in which case the desired output is REFUTES. The reformulator is prompted separately to rewrite the claim into short retrieval text, e.g.
Rewrite the claim as a concise Wikipedia search query. Return only the query text.
The exact prompt strings used in the runs are the ones stored in the released experiment configuration.
Policy training and baselines.
Policies are then trained on frozen all-MiniLM-L6-v2 embeddings with a two-layer MLP and the augmented comp-sum surrogate at $\tau=1$. The L2D, best-fixed, partial-randomization, and joint-random baselines are exactly those defined in Section C.1, instantiated on the FEVER expert pool and advice set.
Reading the main-paper FEVER tables.
The two FEVER tables in Section 5 support distinct claims. Table 1 shows that advice utility is expert-dependent: the best retrieval strategy is not shared across experts, which is precisely the regime in which a joint expert-advice policy is needed. Table 2 then evaluates policies on the true deployment metric, reporting mean and standard deviation over random seeds. Its central message is that our method achieves the lowest mean loss for seven of the eight retained cost regimes and ties L2D when advice becomes too expensive at $\lambda=10$, while also dominating the non-adaptive and randomized references. The best fixed composite action changes with $\lambda$: it is DeBERTa with top-$1$ retrieval for $\lambda\in\{0,0.5,1,2,4\}$, and the cheapest no-advice expert for $\lambda\in\{6,8,10\}$.
Table 7: Best advice action for each expert across the retained FEVER cost grid. Each cell reports the best advice index $k$ and the corresponding true loss. Here $k=0$ denotes no advice, $k=1$ BM25 top-$1$, $k=2$ BM25 top-$3$, $k=3$ BM25 top-$5$, and $k=4$ reformulation followed by BM25 top-$5$.
Expert $\lambda=0$ $\lambda=0.5$ $\lambda=1$ $\lambda=2$ $\lambda=4$ $\lambda=6$ $\lambda=8$ $\lambda=10$
Qwen3-4B-Instruct k=3k{=}3 / 0.667 k=3k{=}3 / 0.682 k=0k{=}0 / 0.695 k=0k{=}0 / 0.695 k=0k{=}0 / 0.695 k=0k{=}0 / 0.695 k=0k{=}0 / 0.695 k=0k{=}0 / 0.695
Qwen3-8B k=3k{=}3 / 0.639 k=3k{=}3 / 0.654 k=3k{=}3 / 0.669 k=3k{=}3 / 0.699 k=0k{=}0 / 0.715 k=0k{=}0 / 0.715 k=0k{=}0 / 0.715 k=0k{=}0 / 0.715
DeBERTa-v3-large k=1k{=}1 / 0.611 k=1k{=}1 / 0.619 k=1k{=}1 / 0.626 k=1k{=}1 / 0.641 k=1k{=}1 / 0.671 k=1k{=}1 / 0.701 k=0k{=}0 / 0.708 k=0k{=}0 / 0.708
Table 7 makes the mechanism explicit over the full retained grid. At low advice cost, different experts prefer different advice actions: the two Qwen experts favor deeper retrieval, whereas DeBERTa prefers the cheapest retrieval action. As $\lambda$ increases, these choices do not collapse simultaneously. Qwen3-4B switches to no advice already at $\lambda=1$, Qwen3-8B remains advice-seeking until $\lambda=4$, and DeBERTa keeps top-$1$ retrieval until $\lambda=6$ before reverting to no advice at $\lambda=8$. This staggered transition is exactly the pattern the joint expert-advice formulation is designed to capture.
Figure 5: Deployed advice distribution of our method at $\lambda=0$ in FEVER. The left panel shows the overall fraction of validation examples assigned to each advice action. The right panel conditions on the routed expert. When advice is free, the learned policy uses queried advice on most examples, but the preferred advice level already varies across experts.
Figure 6: Deployed advice distribution of our method at $\lambda=2$ in FEVER. Compared with Figure 5, the policy shifts mass away from expensive advice actions, but it does not collapse uniformly to no advice. The expert-conditional panel shows that the transition remains expert-dependent.
Figures 5 and 6 show how our method allocates mass across advice actions. As the advice cost increases, the policy shifts away from expensive advice, but it does not do so uniformly. The reallocation toward cheaper actions remains expert-dependent.
Table 8: Expert-by-advice true deferral-advice loss at $\lambda=2$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3 k=4k{=}4
Qwen3-4B-Instruct 0.695 0.726 0.741 0.727 0.710
Qwen3-8B 0.715 0.746 0.749 0.699 0.713
DeBERTa-v3-large 0.708 0.641 0.655 0.677 0.668
Table 9: Expert-by-advice true deferral-advice loss at $\lambda=10$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3 k=4k{=}4
Qwen3-4B-Instruct 0.695 0.846 0.901 0.967 0.790
Qwen3-8B 0.715 0.866 0.909 0.939 0.793
DeBERTa-v3-large 0.708 0.761 0.815 0.917 0.748
Tables 8 and 9 show directly how the deployment-optimal expert-advice pair changes with the cost regime. At $\lambda=2$, the lowest loss is achieved by DeBERTa with top-$1$ retrieval. At $\lambda=10$, the optimum shifts to the cheapest no-query expert. What matters operationally is therefore the full deferral-advice loss, not retrieval quality in isolation.
C.4 Sensitive-Escalation Experimental Details
This appendix reports the tabular sensitive-escalation experiment that is deferred from the main paper. The objective is the same as in FEVER, but in a different regime: advice corresponds to progressively more sensitive feature groups rather than retrieved text. The question is whether the composite expert-advice formulation still improves the true deployment objective once the system must trade predictive value against both expert fees and escalation costs.
Data, action space, and metric.
We use the IEEE-CIS fraud detection benchmark [Howard et al., 2019]. Each example is a payment event with tabular transaction fields, optional identity attributes, and a binary fraud label. A useful operational picture is an online card transaction for which the system initially observes only a restricted low-sensitivity view, such as coarse merchant, amount, and transaction metadata. Before sending the case to an expert, it may decide whether to reveal additional information. For instance, the policy may request contact metadata, a payment-profile summary, a device profile, or stricter identity attributes if these extra signals are worth their escalation cost. We follow the split used in the experiment configuration and train on 120,000120{,}000 examples with an 8,0008{,}000-example validation set. With three experts and five advice actions, the composite action space again has size $3\times 5$. As in FEVER, all policy comparisons follow the global protocol of Section C.1; in particular, the reported numbers are validation averages of the true deferral-advice loss.
Experts and advice construction.
The expert pool contains a logistic-regression model, a histogram GBDT, and an MLP, with consultation costs $\beta=(0.02,0.06,0.09)$. Advice does not take the form of retrieved text here. Instead, it corresponds to revealing additional feature groups beyond the base low-sensitivity transaction view. The advice levels are nested: each action reveals all information from the previous levels and then adds one more group. Concretely, the five advice actions are:
  • $k=0$: no additional information.
  • $k=1$: the base view plus low-security contact metadata.
  • $k=2$: the information in $k=1$ plus medium-security payment-profile attributes.
  • $k=3$: the information in $k=2$ plus high-security device-profile attributes.
  • $k=4$: the information in $k=3$ plus strict identity attributes.
Their base costs are $\gamma^{\mathrm{base}}=(0,0.08,0.12,0.16,0.20)$, multiplied by $\lambda\in\{0,0.04,0.12,0.20,5.0\}$. Richer advice can improve fraud detection, but each escalation level carries a larger operational and privacy cost.
The experts themselves are trained offline in a way that matches this nested advice protocol. For each training example, we do not duplicate the example across all five advice levels. Instead, we sample one advice level uniformly from $\{0,1,2,3,4\}$, reveal exactly the corresponding prefix of feature groups, and train the expert on that masked view augmented with binary reveal indicators. This produces one memory-safe pooled training set in which each expert sees a mixture of low- and high-information cases. The logistic model uses balanced class weights, the histogram GBDT uses class-balanced sample weights, and the MLP additionally oversamples the minority class. After this offline training stage, each expert is evaluated under all five advice levels on the train and validation splits to populate the executed expert-advice cost tensor used by the policy learner.
Policy training and baselines.
Policies are trained on the tabular representation used by the experiment configuration, with the same augmented surrogate at $\tau=1$ as in the FEVER study. For every example and every pair $(j,k)$, we precompute the expert prediction and the resulting realized cost, so all methods are trained and evaluated on exactly the same expert-advice outcome tensor. The augmented policy, L2D, best-fixed, partial-randomization, and joint-random baselines are exactly those defined in Section C.1, instantiated on this fraud-detection expert pool and advice set.
Reading the appendix sensitive-escalation tables.
The tables below play the same role as their FEVER counterparts, but in a different modality. Table 10 first asks how each expert behaves when advice cost changes: which advice action is best for that expert, and how quickly does it revert to no advice as escalation becomes expensive? Table 11 then evaluates full policies on the true deployment metric. The detailed tables at $\lambda=0$, $\lambda=0.12$, and $\lambda=5$ make the cost-dependent expert-advice optimum explicit.
Table 10: Best advice action for each expert across the retained sensitive-escalation cost grid. Each cell reports the best advice index $k$ and the corresponding true loss. Here $k=0$ denotes no advice, $k=1$ low-security metadata, $k=2$ medium-security profile, $k=3$ high-security device profile, and $k=4$ strict identity attributes.
Expert $\lambda=0$ $\lambda=0.04$ $\lambda=0.12$ $\lambda=0.20$ $\lambda=5$
Linear k=4k{=}4 / 0.308 k=3k{=}3 / 0.314 k=0k{=}0 / 0.325 k=0k{=}0 / 0.325 k=0k{=}0 / 0.325
Tree k=4k{=}4 / 0.271 k=4k{=}4 / 0.279 k=2k{=}2 / 0.291 k=2k{=}2 / 0.301 k=0k{=}0 / 0.317
MLP k=2k{=}2 / 0.359 k=2k{=}2 / 0.364 k=0k{=}0 / 0.369 k=0k{=}0 / 0.369 k=0k{=}0 / 0.369
Table 10 highlights the same phenomenon as in FEVER in a tabular setting. The best advice action depends both on the expert and on the cost regime. When escalation is cheap, the linear model prefers the most revealing advice, the tree expert prefers the strict-security tier, and the MLP prefers the medium-security tier. As the escalation cost increases, these preferences shift back toward cheaper advice or no advice, but not at the same rate for all experts. This is precisely the kind of expert-dependent trade-off that motivates a joint expert-advice policy.
Table 11: Validation true deferral-advice loss in the sensitive-escalation experiment. Entries are mean $\pm$ standard deviation over random seeds. Lower is better.
Method $\lambda{=}0$ $\lambda{=}0.04$ $\lambda{=}0.12$ $\lambda{=}0.20$ $\lambda{=}5$
Our 0.260 $\pm$ 0.001 0.267 $\pm$ 0.001 0.274 $\pm$ 0.002 0.279 $\pm$ 0.002 0.281 $\pm$ 0.001
L2D 0.282 $\pm$ 0.001 0.282 $\pm$ 0.001 0.282 $\pm$ 0.001 0.282 $\pm$ 0.001 0.282 $\pm$ 0.001
Best fixed pair 0.271 0.279 0.291 0.301 0.317
Learned expert, random advice 0.287 $\pm$ 0.003 0.297 $\pm$ 0.003 0.304 $\pm$ 0.003 0.313 $\pm$ 0.003 0.837 $\pm$ 0.004
Random expert, learned advice 0.310 $\pm$ 0.003 0.314 $\pm$ 0.003 0.322 $\pm$ 0.003 0.325 $\pm$ 0.004 0.338 $\pm$ 0.003
Random $(j,k)$ 0.325 $\pm$ 0.003 0.330 $\pm$ 0.003 0.339 $\pm$ 0.003 0.348 $\pm$ 0.003 0.885 $\pm$ 0.006
Table 11 shows a consistent pattern across the entire cost sweep. Our method attains the lowest mean true loss at every value of $\lambda$, with the largest gains over L2D when escalation is cheap and only a small residual gap at $\lambda=5$. This is the expected pattern as the optimal policy moves gradually toward no-advice actions. Our method also stays below the best fixed pair throughout the sweep. That fixed pair changes with the cost regime, moving from the tree expert with strict-security advice at $\lambda\in\{0,0.04\}$, to the same expert with medium-security advice at $\lambda\in\{0.12,0.20\}$, and finally to the no-advice tree expert at $\lambda=5$. The partial-randomization ablations further show that the gain comes from coordinating expert selection and advice acquisition rather than from either component alone; in particular, randomizing the advice becomes especially damaging at high cost, because unnecessary escalation is then heavily penalized.
Figure 7: Deployed advice distribution of our method at $\lambda=0.12$. The left panel shows the overall fraction of validation examples assigned to each advice action. The right panel conditions on the routed expert. At this moderate cost, the learned policy still uses queried advice frequently, but the preferred advice level already depends on the chosen expert.
Figure 8: Deployed advice distribution of our method at $\lambda=0.20$. Compared with Figure 7, the policy shifts further toward cheaper advice and no advice. The right panel shows that this transition is still expert-dependent rather than uniform across the routed experts.
Figures 7 and 8 show how the policy allocates mass across advice levels. The transition with $\lambda$ is not simply “query” versus “no-query.” As advice becomes more expensive, the learned policy first moves toward cheaper advice levels and only then collapses toward no advice, which explains why the augmented and L2D risks nearly coincide at $\lambda=5$.
Table 12: Expert-by-advice true deferral-advice loss at $\lambda=0$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3 k=4k{=}4
Linear 0.325 0.322 0.316 0.308 0.308
Tree 0.317 0.301 0.277 0.273 0.271
MLP 0.369 0.382 0.360 0.377 0.372
Table 13: Expert-by-advice true deferral-advice loss at $\lambda=0.12$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3 k=4k{=}4
Linear 0.325 0.332 0.330 0.327 0.332
Tree 0.317 0.310 0.291 0.292 0.295
MLP 0.369 0.391 0.374 0.396 0.396
Table 14: Expert-by-advice true deferral-advice loss at $\lambda=5$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3 k=4k{=}4
Linear 0.325 0.722 0.916 1.108 1.308
Tree 0.317 0.701 0.877 1.073 1.271
MLP 0.369 0.782 0.960 1.177 1.372
Tables 12, 13, and 14 show how the deployment-optimal action changes with the escalation cost. At $\lambda=0$, the tree expert with the strict-security advice tier is best. At $\lambda=0.12$, the optimum has already moved to the medium-security tier for that same expert. At $\lambda=5$, the best action is the no-advice tree expert, and all queried actions are much more costly. The relevant choice is therefore the expert-advice pair with smallest true deferral-advice loss for the current value of $\lambda$.
C.5 CLIP Prompt-Escalation Experimental Details
The setting again matches the deferral-advice protocol, but advice now corresponds to increasingly rich prompt families for frozen CLIP [Radford et al., 2021] experts. The question is whether a joint expert-advice policy improves the true deployment objective when richer prompt sets may improve zero-shot accuracy but also incur an explicit consultation cost.
Data, action space, and metric.
We use ImageNet-1k [Russakovsky et al., 2015]. Each example is an image paired with one of $1000$ class labels. A useful way to read this experiment is as a cost-sensitive zero-shot classification problem: given an image of, say, a goldfish, the system may either classify it with a single canonical prompt such as “a photo of a goldfish” or pay for a richer prompt family before sending the image to a CLIP expert. The retained run uses 39,22239{,}222 training examples and 14,28814{,}288 validation examples. With three experts and four advice actions, the composite action space has size $3\times 4$. As in FEVER and sensitive-escalation, all policy comparisons follow the global protocol of Section C.1; in particular, the reported numbers are validation averages of the true deferral-advice loss.
Experts and advice construction.
The expert pool contains three frozen CLIP models: CLIP-B/32, CLIP-B/16, and CLIP-L/14, with consultation costs $\beta=(0,0.04,0.08)$. Advice does not reveal new image features. Instead, it enlarges the prompt family used to build the text classifier for a given expert. The four advice actions are:
  • $k=0$: no advice, using a single canonical prompt.
  • $k=1$: a clean prompt ensemble with $5$ closely related templates.
  • $k=2$: a medium-diversity family with $16$ templates.
  • $k=3$: a large-diversity family with $32$ templates.
These advice levels are nested in the practical sense that larger prompt families include and extend the simpler ones. Their base costs are $\gamma^{\mathrm{base}}=(0,0.03,0.06,0.09)$, multiplied by $\lambda\in\{0,0.1,0.15,0.3,0.4,0.6,0.7,0.8,1.0,1.5,10.0\}$. This produces a transparent trade-off: richer prompt families can improve zero-shot accuracy, but only at an additional advice cost.
Prompt examples.
It is useful to make these advice actions concrete. For an image whose class is goldfish, the no-advice action uses only the canonical template “a photo of a goldfish.” The clean-ensemble advice augments this with closely related variants such as “a clean photo of a goldfish,” “a centered photo of a goldfish,” and “a high quality photo of a goldfish.” The medium-diversity advice expands further to prompts such as “a close-up photo of a goldfish,” “a bright photo of a goldfish,” “a blurry photo of a goldfish,” and “a black and white photo of a goldfish.” Finally, the large-diversity advice adds still broader views, for instance “a rendering of a goldfish,” “a sketch of a goldfish,” “a nighttime photo of a goldfish,” and “a front-view photo of a goldfish.” In this experiment, acquiring advice therefore means paying to evaluate the image against a richer prompt family before handing it to the selected CLIP expert.
Policy training and baselines.
Policies are trained on frozen CLIP-B/32 image embeddings of dimension $512$, using the same augmented surrogate at $\tau=1$ as in the other experiments. For every image, every expert, and every prompt family, we precompute the expert prediction and the realized cost, so all methods are trained and evaluated on exactly the same expert-advice outcome tensor. The augmented policy, L2D, best-fixed, partial-randomization, and joint-random baselines are exactly those defined in Section C.1, instantiated on this CLIP expert pool and prompt-family advice set.
Reading the appendix CLIP tables.
The tables below mirror those of the previous subsections. Table  15 first shows how each CLIP expert’s preferred advice action changes with the advice cost. Table  16 then evaluates full policies on the true deployment metric, and the detailed tables at $\lambda=0$, $\lambda=0.15$, and $\lambda=10$ make the cost-dependent expert-advice optimum explicit.
Table 15: Best advice action for each expert across the retained CLIP cost grid. Each cell reports the best advice index $k$ and the corresponding true loss. Here $k=0$ denotes the canonical no-advice prompt, $k=1$ the clean ensemble, $k=2$ the medium-diversity prompt family, and $k=3$ the large-diversity prompt family.
Expert $\lambda=0$ $\lambda=0.1$ $\lambda=0.15$ $\lambda=0.3$ $\lambda=0.4$ $\lambda=0.6$ $\lambda=0.7$ $\lambda=0.8$ $\lambda=1$ $\lambda=1.5$ $\lambda=10$
CLIP-B/32 k=3k{=}3 / 0.391 k=2k{=}2 / 0.398 k=2k{=}2 / 0.401 k=0k{=}0 / 0.403 k=0k{=}0 / 0.403 k=0k{=}0 / 0.403 k=0k{=}0 / 0.403 k=0k{=}0 / 0.403 k=0k{=}0 / 0.403 k=0k{=}0 / 0.403 k=0k{=}0 / 0.403
CLIP-B/16 k=3k{=}3 / 0.378 k=2k{=}2 / 0.386 k=2k{=}2 / 0.389 k=0k{=}0 / 0.394 k=0k{=}0 / 0.394 k=0k{=}0 / 0.394 k=0k{=}0 / 0.394 k=0k{=}0 / 0.394 k=0k{=}0 / 0.394 k=0k{=}0 / 0.394 k=0k{=}0 / 0.394
CLIP-L/14 k=3k{=}3 / 0.351 k=1k{=}1 / 0.359 k=1k{=}1 / 0.361 k=1k{=}1 / 0.365 k=0k{=}0 / 0.366 k=0k{=}0 / 0.366 k=0k{=}0 / 0.366 k=0k{=}0 / 0.366 k=0k{=}0 / 0.366 k=0k{=}0 / 0.366 k=0k{=}0 / 0.366
Table 15 shows the same structural pattern as the FEVER and tabular experiments. The best advice action depends on both the expert and the cost regime. When advice is cheap, all three experts prefer queried prompt families, although not the same ones: the two smaller CLIP models move through the medium-diversity family, whereas CLIP-L/14 already switches to the cheaper clean ensemble at $\lambda=0.1$ and $\lambda=0.15$. Once the advice cost increases further, all three experts revert to the no-advice action, but not at the same point. This is again the regime in which a joint expert-advice policy is most natural.
Table 16: Validation true deferral-advice loss in the CLIP prompt-escalation experiment. Entries are mean $\pm$ standard deviation over random seeds. Lower is better.
$\lambda$ Our L2D Best fixed pair Learned expert, random advice Random expert, learned advice Random $(j,k)$
0 0.337 $\pm$ 0.000 0.352 $\pm$ 0.001 0.351 0.343 $\pm$ 0.001 0.374 $\pm$ 0.002 0.379 $\pm$ 0.003
0.1 0.345 $\pm$ 0.000 0.352 $\pm$ 0.001 0.359 0.348 $\pm$ 0.001 0.381 $\pm$ 0.002 0.383 $\pm$ 0.003
0.15 0.348 $\pm$ 0.001 0.352 $\pm$ 0.001 0.361 0.350 $\pm$ 0.001 0.384 $\pm$ 0.002 0.386 $\pm$ 0.003
0.3 0.350 $\pm$ 0.002 0.352 $\pm$ 0.001 0.365 0.357 $\pm$ 0.001 0.387 $\pm$ 0.002 0.392 $\pm$ 0.003
0.4 0.350 $\pm$ 0.002 0.352 $\pm$ 0.001 0.366 0.361 $\pm$ 0.001 0.388 $\pm$ 0.003 0.397 $\pm$ 0.003
0.6 0.350 $\pm$ 0.002 0.352 $\pm$ 0.001 0.366 0.370 $\pm$ 0.001 0.388 $\pm$ 0.002 0.406 $\pm$ 0.003
0.7 0.350 $\pm$ 0.002 0.352 $\pm$ 0.001 0.366 0.374 $\pm$ 0.001 0.388 $\pm$ 0.002 0.410 $\pm$ 0.003
0.8 0.351 $\pm$ 0.002 0.352 $\pm$ 0.001 0.366 0.379 $\pm$ 0.001 0.388 $\pm$ 0.002 0.415 $\pm$ 0.003
1 0.351 $\pm$ 0.001 0.352 $\pm$ 0.001 0.366 0.388 $\pm$ 0.001 0.388 $\pm$ 0.002 0.424 $\pm$ 0.003
1.5 0.351 $\pm$ 0.001 0.352 $\pm$ 0.001 0.366 0.410 $\pm$ 0.001 0.388 $\pm$ 0.002 0.446 $\pm$ 0.003
10 0.352 $\pm$ 0.001 0.352 $\pm$ 0.001 0.366 0.795 $\pm$ 0.003 0.388 $\pm$ 0.002 0.829 $\pm$ 0.004
Table 16 shows a clear cost-dependent transition. Our method attains the lowest mean true loss throughout the low- and moderate-cost regime, improving on both L2D and the best fixed pair. As $\lambda$ increases, the gap narrows, which is expected because the optimal behavior approaches the no-advice regime. At $\lambda=10$, our method is slightly above L2D in mean loss, consistent with the fact that querying has effectively become too expensive to justify. The partial-randomization baselines support the same conclusion as in the other experiments: the gain comes from coordinating expert choice and advice choice rather than from either component in isolation.
Figure 9: Deployed advice distribution of our method at $\lambda=0.3$ in the CLIP prompt-escalation experiment. The left panel shows the overall fraction of validation images assigned to each advice action. The right panel conditions on the routed expert. At this cost level, the policy still queries frequently, but it already favors cheaper prompt families than in the free-advice regime.
Figure 10: Deployed advice distribution of our method at $\lambda=0.4$ in the CLIP prompt-escalation experiment. Compared with Figure 9, the policy already shifts further toward no advice and the cheapest prompt families. The expert-conditional panel shows that this transition remains expert-dependent rather than uniform across routed experts.
Figures 9 and 10 show how the queried mass is distributed across prompt families. As the advice cost increases, the policy moves toward cheaper prompt families and then toward no advice, and it does so differently for different routed experts.
Table 17: Expert-by-advice true deferral-advice loss at $\lambda=0$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3
CLIP-B/32 0.403 0.397 0.392 0.391
CLIP-B/16 0.394 0.385 0.380 0.378
CLIP-L/14 0.366 0.356 0.354 0.351
Table 18: Expert-by-advice true deferral-advice loss at $\lambda=0.15$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3
CLIP-B/32 0.403 0.402 0.401 0.405
CLIP-B/16 0.394 0.390 0.389 0.391
CLIP-L/14 0.366 0.361 0.363 0.364
Table 19: Expert-by-advice true deferral-advice loss at $\lambda=10$. Lower is better.
k=0k{=}0 k=1k{=}1 k=2k{=}2 k=3k{=}3
CLIP-B/32 0.403 0.697 0.992 1.291
CLIP-B/16 0.394 0.685 0.980 1.278
CLIP-L/14 0.366 0.656 0.954 1.251
Tables 17, 18, and 19 show how the deployment-optimal prompt family changes with the advice cost. At $\lambda=0$, all three experts prefer the richest prompt family. At $\lambda=0.15$, the optimum has already shifted to smaller prompt sets. By $\lambda=10$, the no-advice action is uniformly optimal in true loss. The relevant decision is therefore the prompt family with smallest deferral-advice loss, not the richest family available.